### v1

Abstract
Drug repositioning is an economically efficient strategy used to discover new indications for existing drugs beyond their original approvals, expanding their applicability and usage to address challenges in disease treatment. In recent years, deep-learning techniques for drug repositioning have gained much attention. While most deep learning-based research methods focus on encoding drugs and diseases by extracting feature information from neighbors in the network, they often pay little attention to the potential relationships between the features of drugs and diseases, leading to imprecise encoding of drugs and diseases. To address this, we design a dual-feature drug repositioning neural network (DFDRNN) model to achieve precise encoding of drugs and diseases. DFDRNN uses two features to represent drugs and diseases: the similarity feature and the association feature. The model incorporates a self-attention mechanism to design two dual-feature extraction modules for achieving precisely encoding of drugs and diseases: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. The IntraDDFE module extracts features from a single domain (drug or disease domain), while the InterDDFE module extracts features from the mixed domain (drug and disease domain). In particular, the feature is changed by InterDDFE, ensuring a precise encoding of drugs and diseases. Finally, a cross-dual-domain decoder is designed to predict drug-disease associations in both the drug and disease domains. Compared to six state-of-the-art methods, DFDRNN outperforms others on four benchmark datasets, with an average AUROC of 0.946 and an average AUPR of 0.597.
keywords:
Drug repositioning, neural network, dual-feature, association feature, cross-dual-domain decoder
1 Introduction
In today’s Medical field, drug development is complex and time-consuming. Developing a new drug involves three stages: the discovery stage, the preclinical stage, and the clinical stage, which typically takes 10-15 years, with an average cost of up to \$2.8 billion fisher2006alterations ; vijayan2022enhancing . Although scientists continue to discover new drug molecules through unremitting efforts and innovation, the traditional drug development model often faces many risks, including the termination of drug development due to drug efficacy, drug safety, or commercial concerns. Remarkably, a significant proportion (80-90%) of these projects fail in clinical trials vijayan2022enhancing ; dimasi2010trends . Therefore, finding a more efficient and precise method for drug development has become critically important.
Drug repositioning, a strategy that garnered significant attention, is proposed to explore alternative indications and potential side effects for already-licensed medications dudley2011exploiting ; wang2024landscape . Compared with traditional drug research and development methods, this approach offers several advantages, such as shorter time, lower cost, and reduced risk. Computational drug repositioning, a favored method among researchers, can significantly reduce the time and energy required to identify candidate drugs. Computational drug repositioning methods are divided into four categories cai2023machine ; luo2021biomedical : network diffusion-based methods, machine learning-based methods, matrix factorization-based methods, and deep learning-based methods.
Network diffusion-based methods classify the drug reposition-ing problem as a link prediction issue within graph neural networks kumar2020link . These methods focus on building heterogeneous networks containing multiple types of biological entities and linking drugs to diseases through message propagation on paths bridging different networks. A commonly used method is a random walk and its variants luo2021biomedical ; lagcn . For instance, Martinez et al. martinez2015drugnet proposed a network-based drug repositioning prioritization method called DrugNet. It integrates information about diseases, drugs, and proteins to simultaneously prioritize drug-disease and disease-drug associations. Another example is Luo et al.’s method BIRW_Cdataset , which effectively predicted potential associations between drugs and diseases by constructing an integrated similarity network between drugs and diseases and combining it with a bidirectional random walk (BiRW).
Machine learning-based methods classify drug repositioning as a binary classification problem. These methods typically use traditional machine learning techniques for drug-disease association prediction, including random forest (RF), support vector machine (SVM), and decision trees cai2023machine . For example, Peng et al. screening designed a novel DTI screening framework called PUDTI. PUDTI utilizes SVM to infer candidate targets for drug repositioning. Shi et al. shi2019lasso proposed a method for predicting drug-target interactions based on Lasso dimensionality reduction and a random forest classifier. Zhao et al. HINGRL proposed HINGRL, which uses a random forest classifier in the final decoding step to predict drug-disease association. In addition, there are decision trees gradient , linear regression, and logistic regression you2019predicting ; elisei2019recursive .
Matrix decomposition-based methods categorize the drug repositioning problem as a recommendation system issue. These methods employ matrix factorization techniques to reduce dimensionality and extract key features, thus better capturing the associations between drugs and diseases. Some standard matrix factorization techniques include singular value decomposition (SVD) and non-negative matrix factorization (NMF). For example, Yang et al. MSBMF proposed a Multiple Similarity Bilinear Matrix Factorization (MSBMF) method, where the similarity matrices of drugs and diseases are concatenated and the drug-disease association matrix is reshaped through matrix decomposition. Luo et al. luo2018SVT employed a fast Singular Value Thresholding (SVT) algorithm to propose a drug repositioning recommendation system (DRRS). Additionally, Sadegh et al. sadeghi2022NMF-DR and Xuan et al. xuan2019dd combined non-negative matrix factorization to predict drug-disease associations.
In recent years, methods based on deep learning have become a hot research topic for drug repositioning. They can effectively capture high-level hidden representations from various data sources and demonstrate state-of-the-art performance. Unlike traditional machine learning methods, deep learning approaches can learn more abstract information by building deep architectures without manual feature selection and tuning yu2022deep . Generally, methods based on deep learning for drug repositioning often integrate graph neural networks such as Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to construct models. For example, Yu et al. lagcn integrated a drug-disease heterogeneous network with a GCN network to build a multi-layer connected end-to-end network model called LAGCN. They also incorporated layer attention mechanisms for drug-disease association prediction. Gu et al. gu2022REDDA designed a biological network of five biological entities and equipped it with three attention mechanisms to predict drug-disease associations. Wang et al. wang2023MSGCA designed a multi-similarity graph convolutional autoencoder(MSGCA). Sun et al. sun2022PSGCN predicted new drug-disease associations by constructing partner-specific graphs for drug-disease pairs and combining them with the GCN. By incorporating the GCN, Xuan et al. xuan2022heterogeneous constructed multiple drug-disease heterogeneous networks based on different drug similarities and encoded drugs and diseases at Multi-scale neighbors. Cai et al. cai2021DRHGCN designed intra-domain and inter-domain feature extraction modules through the GCN and the bilinear aggregator, proposing DRHGCN. Meng et al. DRAGNN utilized a GAT for encoding drugs and diseases and introduced a weighted local information augmented graph neural network model named DRAGNN. Most current deep learning-based methods for drug repositioning typically utilize the similarity between neighbors and aggregate neighbor features for training, achieving considerable results. However, many end-to-end deep learning models do not involve purposeful encoding. They delegate the task to the optimizer, relying solely on the coordinated training of the encoding and decoding processes to complete association prediction. This may lead to the introduction of a large amount of noisy information during model training, resulting in imprecise encoding of drugs and diseases and affecting the prediction results.
This paper proposes a dual-feature drug repositioning neural network (DFDRNN) model to address the above issue. The DFDRNN constructs two features for drugs and diseases: the similarity feature and the association feature. It utilizes the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module to extract drug and disease features and achieve precise encoding. A self-attention mechanism (SAM) is employed in these two modules to capture the complex relationships between adjacent nodes. Finally, the decoder adopts a cross-dual-domain decoder to achieve precise prediction of drug-disease associations. Compared to GCN-based methods, DFDRNN employs a SAM to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation. This enables efficient aggregation of neighbors and enhances the model’s predictive power. We compare DFDRNN with six state-of-the-art methods on four benchmark datasets. Experimental results show that our model outperforms other models. This article makes the following contributions:
(1) We propose a dual-feature method to represent drugs and diseases. The proposed approach involves two features: the similarity feature and the association feature. These two features are independent of each other in the network, while they can be extracted simultaneously to represent drugs and diseases in each network layer.
(2) Two dual-feature extraction modules are proposed: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. In particular, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature. The proposed two modules effectively capture the features of drugs and diseases, ensuring a precise encoding of drugs and diseases.
(3) A cross-dual-domain decoder is proposed, implementing a cross-computation between two encoded features to obtain decoding results for the drug and disease domains. The fusion of the two decoding results improves the accuracy of association prediction.
The remainder of this paper is organized as follows. Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results; and Section 4 concludes our work.
2 Materials and Methods
2.1 Datasets
We conducted experiments on four benchmark datasets. The first dataset, Fdataset Fdataset , includes 593 drugs from DrugBank knox2024drugbank and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) hamosh2005OMIM database, comprising 1933 drug-disease associations. The second dataset, Cdataset BIRW_Cdataset , consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations. The third dataset, Ldataset lagcn , contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) davis2023CTD . The fourth dataset, LRSSL liang2017lrssl , comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations. Table 1 summarizes the four datasets’ statistics.
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation weininger1988smiles . The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit steinbeck2003chemistry . Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner van2006text . MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database lipscomb2000medical .
Table 1: The statistics of the four datasets used in this study
[…]
Cdataset 663 409 2532
Ldataset 269 598 18416
LRSSL 763 681 3051
[…]
2.2 Model structure
Our research is focused on predicting potential associations between drugs and diseases. To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model. The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the feature. A self-attention (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes. The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains. Figure 1 illustrates the diagram of the proposed DFDRNN model.
Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks. The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on self-attention. A layer attention mechanism is also introduced to compute the final similarity and association features. Finally, the decoder calculates drug-disease association scores by decoding separately for the drug and disease domains.
2.2.1 Three networks
To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity of the two drug incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge. We use three matrices to represent these three networks.
Specifically, suppose that there are $n$ drugs (denoted by $r_{1},r_{2},\ldots,r_{n}$) and $m$ diseases (denoted by $d_{1},d_{2},\ldots,d_{m}$). We use $S^{r}$ (or $S^{d}$) to denote the adjacency matrix of a drug-drug similarity network (or disease-disease similarity network), which is a $n\times n$ ($m\times m$) real matrix such that the entry $s_{ij}\in[0,1]$ denotes the degree of similarity of drug $i$ and drug $j$ (or disease $i$ and disease $j$), where $i,j\in\{1,2,\ldots,n\}$ (or $i,j\in\{1,2,\ldots,m\}$). For a drug-disease association network, we use a $n\times m$ binary matrix, denoted by $A$, to represent it, where the entry $a_{ij}=1$ if drug $i$ is associated with disease $j$ and $a_{ij}=0$ if drug $i$ is not associated with disease $j$ or the associativity is yet to be observed, for $i=1,2,\ldots,n$ and $j=1,2,\ldots,m$.
To reduce the impact of data noise, we aggregate only the top-$t$ neighbors for each drug in $S^{r}$ and each disease in $S^{d}$, and then two new matrices $R\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{m\times m}$ are generated to replace $S^{r}$ and $S^{d}$, respectively, where the entries $r_{ij}$ ($i,j\in\{1,2,\ldots,n\}$) of $R$ and $d_{ij}$ ($i,j\in\{1,2,\ldots,m\}$) of $D$ are defined as follows. For a drug $r_{i}$ (and a disease $d_{i}$), we denote by $N_{k}(r_{i})$ (and $N_{k}(d_{i})$) the set of the $t$-nearest neighbors of $r_{i}$ in $S^{r}$ (and $d_{i}$ in $S^{d}$), i.e., the weight of edge $r_{i}x$ (and edge $d_{i}y$) for every $x\in N_{k}(r_{i})$ (and every $y\in N_{k}(d_{i})$) is top-$t$ among all neighbors of $r_{i}$ (and $d_{i}$), and let $\tilde{N_{k}}(r_{i})=N_{k}(r_{i})\cup\{r_{i}\}$ (and $\tilde{N_{k}}(d_{i})=N_{k}(d_{i})\cup\{d_{i}\}$) be the extended $t$-nearest neighbors of $r_{i}$ (and $d_{i}$).
$\begin{matrix}R_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{k}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&D_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{k}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)
Finally, to encode drugs and diseases simultaneously and distinguish IntraDDFE from InterDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$. $E_{s}$ is used for IntraDDFE, while $E_{a}$ is used for InterDDFE, defined as:
$\ \qquad\qquad E_{s}=\begin{bmatrix}R&0\\ 0&D\end{bmatrix}\qquad E_{a}=\begin{bmatrix}0&A\\ A^{T}&0\end{bmatrix}$ (2)
2.2.2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q$ and a group of $K$-$V$ pairs to a new output vector. The output of the attention is a weighted sum of $V$ vectors, where the weight for each $V$ is computed by taking the dot product of the $Q$ with the corresponding $K$ vaswani2017attention . In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation sam to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of sam can be described as follows:
$Q=H^{in}w^{Q},\qquad K=H^{in}w^{K},\qquad V=H^{in}w^{V}$ (3)
where $w^{Q},w^{K},w^{V}\in\mathbb{R}^{\mathit{k}\times k}$ are three trainable $k\times k$ real matrices. Clearly, $Q,K,$ and $V$ are $(n+m)\times k$ real matrices.
Each $Q$, $K$, and $V$ is partitioned into $p$ parts (called heads), denoted by $Q^{i}$,$K^{i}$, and $V^{i}$, $i=1,2,\ldots,p$, and the dimension of each part (i.e., the number of columns) is $k^{\prime}=k/p$. We use $Q^{q}_{i}$, $K^{q}_{i}$, and $V^{q}_{i}$ to denote the feature of node $i$ in $Q^{q}$, $K^{q}$, and $V^{q}$ (i.e., the row vector corresponding to $i$), respectively. Let $X^{q}$ be the $q$-th head of $X\in\{Q,K,V\}$. Then, the weight of aggregating neighbor information is calculated based on $Q^{q}$ and $K^{q}$. Specifically, for a node $i$ (drug or disease), the weight $\beta^{r}_{ij}$ of $i$ aggregating its neighbor $j$ is defined as
$\beta^{q}_{ij}=\frac{exp(\frac{Q^{q}_{i}\bigodot K^{q}_{j}}{\sqrt{k^{\prime}}})}{\sum_{z\in N(i)}exp(\frac{Q^{q}_{i}\bigodot K^{q}_{z})}{\sqrt{k^{\prime}}})}$ (4)
[…]
$H_{\texttt{sam}}^{i}=\texttt{concat}(\sum_{j\in N(i)}\beta^{1}_{ij}V^{1}_{j},\sum_{j\in N(i)}\beta^{2}_{ij}V^{2}_{j},\ldots,\sum_{j\in N(i)}\beta^{p}_{ij}V^{p}_{j})$ (5)
Moreover, let $H_{\texttt{sam}}$ represent the feature matrix obtained by SAM, $H_{r}\in\mathbb{R}^{n\times k}$ be the drug feature obtained by SAM, and $H_{d}\in\mathbb{R}^{m\times k}$ is the disease feature obtained by SAM. Then, we have
$H_{r}=\begin{bmatrix}H^{r_{1}}_{\texttt{sam}}\\ H^{r_{2}}_{\texttt{sam}}\\ \dots\\ H^{r_{m}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\texttt{sam}}\\ H^{d_{2}}_{\texttt{sam}}\\ \dots\\ H^{d_{n}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{\texttt{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
A fully connected layer is applied after the SAM function to extract further features of drugs and diseases. We use the notation samf to denote the whole process, i.e., samf is the combination of sam with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU maas2013leakyrelu activation function. Then,
$\texttt{samf}(E,H^{in})$ $=\texttt{FC}(\texttt{sam}(E,H^{in}))$ (7)
$=\begin{bmatrix}\sigma(H_{r}w_{r}+b_{r})\\ \sigma(H_{d}w_{d}+b_{d})\end{bmatrix}$
2.2.3 Encoder
The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: IntraDDFE module based on network $E_{s}$ and InterDDFE module based on network $E_{a}$. First, the initial features of drugs and diseases are generated in two features, including the similarity feature $H^{init}_{s}$ and the association feature $H^{init}_{a}$, described as follows.
$\ \qquad\qquad H^{init}_{s}=\begin{bmatrix}S^{r}&0\\ 0&S^{d}\end{bmatrix}\qquad H^{init}_{a}=\begin{bmatrix}0&A\\ A^{T}&0\end{bmatrix}$ (8)
Next, DFDRNN projects the initial features to a $k$-dimensional space through a linear layer. The similarity feature $H_{s}^{0}$ and the association feature $H_{a}^{0}$ in the 00-layer are defined as follows:
$H^{0}_{s}=H^{init}_{s}\times M\quad,\quad H^{0}_{a}=H^{init}_{a}\times M$ (9)
where $M\in\mathbb{R}^{(n+m)\times k}$ is trainable parameter matrix.
Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$. In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\texttt{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE. Then,
$(\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a})$ $=f^{\ell}(E_{s},H^{\ell}_{s},H^{\ell}_{a})$ (10)
$=(\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
For the InterDDFE module, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer. Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\texttt{samf}^{\ell}_{g}$ represents the aggregation function. Then,
$(\tilde{H}^{\ell+1}_{a},\tilde{H}^{\ell+1}_{s})$ $={g^{\ell}(E_{a},H^{\ell}_{s},H^{\ell}_{a})}$ (11)
${=(\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
It should be noted that compared with the IntraDDFE module, the InterDDFE module changes the features: the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$ and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
Finally, for each module, the similarity (and association) features (including one input feature and two output features corresponding to the two modules) in the $\ell$-layer are added into an input similarity (and association) feature of the $(\ell+1)$-layer, i.e.,
$H^{\ell+1}_{s}=\hat{H}^{\ell+1}_{s}+\tilde{H}^{\ell+1}_{s}+H^{\ell}_{s}$ (12)
$H^{\ell+1}_{a}=\hat{H}^{\ell+1}_{a}+\tilde{H}^{\ell+1}_{a}+H^{\ell}_{a}$ (13)
As the depth of network layers increases, the valuable information for predicting drug-disease association extracted by the DFDRNN model decreases. Therefore, we introduce a layer attention mechanism to allocate different weights to the results of different layers. We use $L$ to represent the total number of layers in DFDRNN. The weight of the $\ell$-th layer, denoted by $\beta^{\ell}$, is initialized by $1/L$ and automatically learned by the model. Then, the final encoding of the similarity feature and the association feature $H^{r}_{s}$, $H^{r}_{a}$, $H^{d}_{s}$, and $H^{d}_{a}$ of drugs and diseases are obtained by implementing the layer attention mechanism as follows.
$\begin{bmatrix}H^{r}_{s}\\ H^{d}_{s}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{s}),\qquad\begin{bmatrix}H^{r}_{a}\\ H^{d}_{a}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{a})$ (14)
2.2.4 Decoder
A cross-dual-domain decoder is utilized to predict drug-disease association. We perform cross computation between two features, using $s_{r}\in\mathbb{R}^{n\times m}$ and $s_{d}\in\mathbb{R}^{n\times m}$ to represent the prediction values for the drug domain and the disease domain, respectively. Let $s\in\mathbb{R}^{n\times m}$ denote the final score matrix of drug-disease associations, where $s_{ij}$ represents the prediction score between drug $r_{i}$ and disease $d_{j}$. That is,
${s_{r}}$ $=\texttt{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
${s_{d}}$ $=\texttt{sigmod}(H^{d}_{a}\times(H^{r}_{s})^{T})$ (16)
${s}$ $={(s_{r}+(s_{d})^{T})/2.0}$ (17)
Algorithm 1 gives a detailed explanation of predicting novel drug-disease associations using DFDRNN.
Algorithm 1 DFDRNN
1:Input: Two heterogeneous network ($E_{s}$ and $E_{a}$); initial the similarity feature and the association feature (${H}^{init}_{s}$ and ${H}^{init}_{a}$); number of layers $L$; maximum training epochs $C$.
2:Output: Drug-disease association score matrix $s$.
3:Initialize the weight matrix ($w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$).
4:for $c=1$ to $C$ do
5:     Compute the similarity feature $H^{0}_{s}$ and the association feature $H^{0}_{a}$ with Equation (9).
6:     for $i=1$ to $L$ do
7:         Compute IntraDDFE $\hat{H}^{\ell}_{s},\hat{H}^{\ell}_{a}$ with Equation (10).
8:         Compute InterDDFE $\tilde{H}^{\ell}_{s},\tilde{H}^{\ell}_{a}$ with Equation (11).
9:         Fuse the same feature and the output of the previous layer yields $H^{\ell}_{s}$ and $H^{\ell}_{a}$ with Equation (12) (13)
10:     end for
11:     Utilizing layer attention mechanism to compute the final embedding with Equation (14)
12:     Obtain Computer the prediction score matrix $s$ with Equation (15)(16)(17)
13:     Update the parameters $w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$ with Equation (18).
14:     $c\leftarrow c+1$.
15:end for
2.3 Optimization
Predicting the drug-disease association can be viewed as a binary classification task. The objective is to determine whether a given drug is associated with a specific disease. Given the serious imbalance between the number of positive samples (drug-disease associations) and negative samples (drug-disease non-associations), we employ binary cross-entropy as the loss function. We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively. We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($i$)-disease($j$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
$loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{i,j})+\sum_{(i,j)\in y^{-}}\log(1-s_{i,j}))$ (18)
We optimize the DFDRNN model using the Adam kingma2014adam optimizer and initialize the parameters’ weights with Xavier initialization glorot2010understanding . We randomly drop some edges and apply regular dropout srivastava2014dropout to improve the model’s generalization ability.
3 Experiments and Results
3.1 Baseline methods
To assess DFDRNN’s performance, we compared it with six state-of-the-art drug-disease association prediction models across the above four datasets. These models are:
$\bullet$ LAGCN lagcn : This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN cai2021DRHGCN : This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF meng2022DRWBNCF : This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA zhang2024NCH-DDA : This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT huang2024HDGAT : This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN tang2023DRGBCN : This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
3.2 Experimental parameter configuration
The model’s parameter values are as follows: DFDRNN’s feature embedding dimension is 128; the number of network layers is 3; the number of attention heads is p = 2; the dropout rate is 0.4; the edge dropout rate is 0.5 on the Ldataset and 0.2 on other datasets; the learning rate is 0.008; and the maximum training epoch is 800. Additionally, we set top-$t$ to 7 based on parameter sensitivity experiments.
3.3 Evaluation metrics
Table 2: The statistics of the four datasets used in this study
Actual Predicted Positive Negative
Positive TP FN
Negative FP TN
We categorize the prediction results into four sets. As shown in Table 2, TP represents the number of true positives correctly classified as positive, FN represents the number of false negatives incorrectly classified as negative, FP represents the number of false positives incorrectly classified as positive, and TN represents the number of true negatives correctly classified as negative. We then calculate True Positive Rate (TPR or Recall), False Positive Rate (FPR), and precision as follows:
[…]
$\text{FPR}=\frac{FP}{FP+TN}$ (20)
$\text{Precision}=\frac{TP}{TP+FP}$ (21)
Calculating the FPR, TPR, and precision at various thresholds, we construct the receiver operating characteristic (ROC) curve, with FPR on the X-axis and TPR on the Y-axis. Additionally, we generate the precision-recall (PR) curve, with recall on the X-axis and precision on the Y-axis. The evaluation metrics used for DFDRNN are the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
3.4 Performance of DFDRNN in the cross-validation
We employ the 10-fold cross-validation method to assess the model’s performance. During this process, the dataset is randomly partitioned into ten mutually exclusive subsets, each containing 10% known drug-disease association pairs and 10% unknown pairs. One subset serves as the test set, while the remaining nine subsets are used to train the model. This process is repeated ten times (i.e., 10-fold cross-validation) with different random seeds, and the average and standard deviation of the results are calculated to compare the model’s performance. The experimental results, including AUROC and AUPR values, are presented in Table 3. The ROC and PR curves can also be found in Supplementary Figures S1-S4.
In Table 3, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets. This performance surpassed the second-ranked NCH-DDA model by 0.75% and 3.1% in AUROC and AUPR, respectively. DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations. It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
Table 3: The AUROCs and AUPRs of DFDRNN and the other comparative models under 10-times 10-fold cross-validation.
[…]
AUROC LAGCN 0.881±0.002 0.911±0.002 0.864±0.001 0.935±0.001 0.898
DRHGCN 0.949±0.001 0.965±0.001 0.864±0.001 0.960±0.001 0.935
DRWBNCF 0.924±0.002 0.942±0.001 0.821±0.001 0.935±0.001 0.906
NCH-DDA 0.954±0.002 0.959±0.002 0.879±0.001 0.963±0.002 0.939
HDGAT 0.917±0.002 0.939±0.001 0.874±0.001 0.936±0.001 0.916
DRGBCN 0.931±0.003 0.946±0.001 0.824±0.002 0.943±0.001 0.911
DFDRNN 0.960±0.001 0.973±0.001 0.885±0.001 0.965±0.001 0.946
AUPR LAGCN 0.197±0.005 0.307±0.008 0.522±0.002 0.254±0.002 0.320
DRHGCN 0.531±0.003 0.619±0.003 0.523±0.001 0.408±0.002 0.520
DRWBNCF 0.493±0.001 0.569±0.004 0.413±0.003 0.349±0.005 0.456
NCH-DDA 0.623±0.006 0.675±0.003 0.552±0.002 0.466±0.002 0.579
HDGAT 0.505±0.005 0.566±0.008 0.546±0.003 0.403±0.005 0.505
DRGBCN 0.408±0.007 0.451±0.006 0.434±0.005 0.262±0.006 0.388
DFDRNN 0.623±0.003 0.688±0.002 0.573±0.001 0.502±0.002 0.597
[…]
{tablenotes}
The best results in each row are in bold faces and the second best results are underlined.
3.5 Effect of the number of nearest neighbors
Figure 2: (a)Line plot of AUROC as a function of top-$t$ on four benchmark datasets; (b)Line plot of AUPR as a function of top-$t$ on four benchmark datasets; (c)Line plot of the average AUROC as a function of top-$t$ on four datasets; (d)Line plot of the average AUPR as a function of top-$t$ on four datasets.
In the context of DFDRNN, the value of top-$t$ determines the number of neighbors features aggregated during IntraDDFE, which influences the model’s performance. A small value of $t$ limits the neighbor information obtained, thereby affecting the model’s performance. On the other hand, as $t$ increases, more neighbors are obtained, but this also introduces noise information, leading to the propagation of unnecessary data. The performance of DFDRNN under 10-fold cross-validation varies with $t$ on the four datasets, as demonstrated in Figure 2 for $t\in\{1,2,3,\ldots,20\}$.
Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$. However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention aggregation capability. Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across four datasets. For detailed numerical values, please refer to Supplementary Tables S1 and S2.
3.6 Ablation analysis
3.6.1 Encoder
The encoder of DFDRNN incorporates SAM to encode drugs (or diseases) into the similarity feature (SF) and the association feature (AF). We introduce two dual-feature extraction modules, IntraDDFE and InterDDFE, to precisely encode drugs and diseases, with InterDDFE changing the feature (CF) of drugs (or diseases). We developed various model variants to underscore these elements’ importance in the DFDRNN encoder and evaluated their performance.
  • DFDRNN-SF: Remove the association feature and encode drugs and diseases solely as the similarity feature.
  • DFDRNN-AF: Remove the similarity feature and encode drugs and diseases solely as the association feature.
  • DFDRNN-GCN: SAM was replaced by a GCN to observe the impact on the model’s performance.
  • DFDRNN-noCF: Remove the change of features in InterDDFE, meaning that the similarity feature and the association feature will not be changed into each other after InterDDFE.
Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-fold cross-validation. By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Fdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets. However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature. Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance. Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information. Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Fdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the InterDDFE module.
Table 4: The AUROCs and AUPRs of DFDRNN and its four variants under 10-fold cross-validation
[…]
AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
DFDRNN-SF 0.958 0.612 0.972 0.690 0.873 0.538 0.964 0.464
DFDRNN-AF 0.942 0.606 0.962 0.678 0.884 0.570 0.950 0.504
DFDRNN-GCN 0.950 0.620 0.964 0.671 0.874 0.551 0.956 0.483
DFDRNN-noCF 0.959 0.611 0.970 0.676 0.868 0.523 0.963 0.462
DFDRNN 0.960 0.622 0.973 0.689 0.885 0.572 0.965 0.504
[…]
3.6.2 Decoder
Figure 3: non-cross dual-domain decoder
To demonstrate the effectiveness of the cross-dual-domain decoder, an experiment was designed to compare the AUROC value and the AUPR value obtained by DFDRNN with those obtained by five variants of DFDRNN, labeled as $s_{r},s_{d},s^{non},s^{non}_{r}$, and $s^{non}_{d}$. Here, $s_{r}$ (Equation 15), $s_{d}$ (Equation 16), and $s^{non}$ represent the models derived from DFDRNN by substituting the cross-dual-domain decoder with a single drug-domain decoder, a single disease-domain decoder and a non-cross dual-domain decoder encompassing decoding results of both the drug and disease domains (as shown in Figure 3), respectively. Similarly, $s^{non}_{r}$ and $s^{non}_{d}$ are models obtained from $s^{non}$ by replacing the non-cross dual-domain decoder with a single drug-domain decoder and a single disease-domain decoder, respectively. The specific implementation of the non-cross dual-domain decoder is described as follows.
${s^{non}_{r}}$ $=\texttt{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
${s^{non}_{d}}$ $=\texttt{sigmod}(H^{d}_{s}\times(H^{r}_{s})^{T})$ (23)
${s^{non}}$ $={(s^{non}_{r}+(^{non}_{d})^{T})/2.0}$ (24)
The heatmaps, a crucial visual representation of the results, are shown in Figure 4 across the four datasets: Fdataset, Cdataset, Ldataset, and LRSSL. DFDRNN’s consistent outperformance in achieving the highest AUROC and AUPR values across all four datasets underscores the significance of the cross-dual-domain decoder. Furthermore, the performance of $s_{r}$, $s_{d}$, $s^{non}_{r}$, and $s^{non}_{d}$ consistently lags behind that of DFDRNN and $s^{non}$, further highlighting the effectiveness of dual-domain decoding.
Figure 4: Performance of different decoding methods: (a) AUROC heatmap; (b) AUPR heatmap.
3.7 Identification of candidates for new diseases
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic singh2020drug . Finding suitable drugs for emerging diseases is a complex and crucial task. To evaluate DFDRNN’s predictive ability in identifying potential candidate drugs for new diseases, we conducted leave-one-out cross-validation (LOOCV) on the Fdataset dataset. Specifically, for each disease $d_{x}$, all drug-disease associations related to $d_{x}$ were excluded from the analysis and treated as test samples. The remaining drug-disease associations were used as training samples, relying solely on the similarity between disease $d_{x}$ and other diseases to predict its candidate drugs.
Figure 5: The performance of all methods in predicting potential drugs for new diseases on Fdataset: (a) Receiver operating characteristic (ROC) curves of prediction results obtained by applying DFDRNN and other competitive methods; (b) precision-recall (PR) curves of prediction results obtained by applying DFDRNN and other competitive methods.
In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated. The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCHDDA by 0.99%. Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
3.8 Case study
To assess the real-world performance of DFDRNN, we conducted case studies on two diseases (Alzheimer’s disease (AD) and Parkinson’s disease (PD)) from the Fdataset. We utilized all drug-disease associations in the Fdataset as the training set to train DFDRNN and predict the association probabilities for all unknown drug-disease pairs. Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD. To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB knox2024drugbank , PubChem kim2016pubchem , CTD davis2023CTD , DrugCentral avram2021drugcentral , and ClinicalTrials.
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table 5 displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment erkulwater1989amantadine . Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients lemke1995effect . DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
Table 5: The top 10 DFDRNN-predicted candidate drugs for AD
[…]
2 DB00564 Carbamazepine CTD
3 DB04844 Tetrabenazine CTD
4 DB00682 Warfarin CTD
5 DB01241 Gemfibrozil CTD/PubChem/ ClinicalTrials
6 DB00160 Alanine CTD/ClinicalTrials
7 DB00903 Etacrynic acid CTD
8 DB00928 Azacitidine CTD
9 DB00482 Celecoxib CTD
10 DB01262 Decitabine CTD
[…]
Table 6: The top 10 DFDRNN-predicted candidate drugs for PD
[…]
2 DB01068 Clonazepam CTD/ClinicalTrials
3 DB01235 Levodopa DB/CTD/PubChem/ DrugCentral/ClinicalTrials/
4 DB00810 Biperiden DB/CTD/PubChem/ DrugCentral
5 DB00424 Hyoscyamine DB
6 DB00575 Clonidine ClinicalTrials
7 DB00915 Amantadine DB/PubChem/DrugCentral/ ClinicalTrials
8 DB00989 Rivastigmine DB/CTD/PubChem/ DrugCentral/ClinicalTrials
9 DB00376 Trihexyphenidyl DB/CTD/PubChem/ DrugCentral
10 DB00215 Citalopram CTD
[…]
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provide reassurance about the reliability of the model.
4 Conclusion
Our study introduces DFDRNN, a neural network model designed for drug repositioning. DFDRNN uses IntraDDFE and InterDDFE modules to extract the similarity and association features of drugs (or diseases), respectively, achieving precise encoding of drugs (or diseases). Compared with the IntraDDFE module, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature. In both modules, we use the self-attention mechanism to aggregate the neighbor information of drugs (and diseases). Additionally, the DFDRNN model fuses three similarity (or association) features into a new similarity (or association) feature and uses a layer attention mechanism to calculate the final similarity (or association) feature. Finally, a cross-dual-domain decoder process is adopted to predict drug-disease associations in both the drug and disease domains. DFDRNN performs best in 10-fold cross-validation across four benchmark datasets and achieves the highest AUROC in leave-one-out cross-validation. The experimental results confirm the effectiveness of our proposed DFDRNN model in drug repositioning.
In the future, we will consider incorporating more biological entities, such as miRNA and proteins, for drug repositioning. Additionally, we will explore more accurate methods to calculate drug (or disease) similarity to reduce data noise.
Key Points We propose DFDRNN, a new deep learning method for drug repositioning, which simultaneously extracts two kinds of features: the similarity feature and the association feature to represent drugs and diseases in each network layer. Based on two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module, DFDRNN captures the potential relationships between the features of drugs and diseases. A cross-dual-domain decoder is proposed to implement a cross-computation between two encoded features to redict drug-disease associations. DFDRNN outperformed the state-of-the-art methods on four benchmark datasets.
5 Acknowledgments
This work was supported by the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China under grants 62172072 and 62272115.

### v2

@@
-
+Conceptualization of this study, Methodology, Software
 Abstract
-Drug repositioning is an economically efficient strategy used to discover new indications for existing drugs beyond their original approvals, expanding their applicability and usage to address challenges in disease treatment.
-In recent years, deep-learning techniques for drug repositioning have gained much attention.
-While most deep learning-based research methods focus on encoding drugs and diseases by extracting feature information from neighbors in the network, they often pay little attention to the potential relationships between the features of drugs and diseases, leading to imprecise encoding of drugs and diseases.
-To address this, we design a dual-feature drug repositioning neural network (DFDRNN) model to achieve precise encoding of drugs and diseases.
-DFDRNN uses two features to represent drugs and diseases: the similarity feature and the association feature.
-The model incorporates a self-attention mechanism to design two dual-feature extraction modules for achieving precisely encoding of drugs and diseases: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
-The IntraDDFE module extracts features from a single domain (drug or disease domain), while the InterDDFE module extracts features from the mixed domain (drug and disease domain).
-In particular, the feature is changed by InterDDFE, ensuring a precise encoding of drugs and diseases.
-Finally, a cross-dual-domain decoder is designed to predict drug-disease associations in both the drug and disease domains.
-Compared to six state-of-the-art methods, DFDRNN outperforms others on four benchmark datasets, with an average AUROC of 0.946 and an average AUPR of 0.597.
+Uncovering new therapeutic uses of existing drugs, drug repositioning offers a fast and cost-effective strategy and holds considerable significance in the realm of drug discovery and development.
+In recent years, deep learning techniques have emerged as powerful tools in drug repositioning due to their ability to analyze large and complex datasets.
+However, many existing methods focus on extracting feature information from nearby nodes in the network to represent drugs and diseases, without considering the potential inter-relationships between the features of drugs and diseases, which may lead to inaccurate representations.
+To address this limitation, we use two features (similarity and association) to capture the potential relationships between the features of drugs and diseases, proposing a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model.
+DFDRNN uses a self-attention mechanism to extract neighbor features and incorporates two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module for extracting features within a single domain (drugs or diseases) and the inter-domain dual-feature extraction (InterDDFE) module for extracting features across domains.
+By utilizing these modules, we ensure more appropriate encoding of drugs and diseases.
+Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains.
+Our proposed DFDRNN model outperforms six state-of-the-art methods on four benchmark datasets, achieving an average AUROC of 0.946 and an average AUPR of 0.597.
+Case studies on two diseases show that the proposed DFDRNN model can be applied in real-world scenarios, demonstrating its significant potential in drug repositioning.
 keywords:
-Drug repositioning, neural network, dual-feature, association feature, cross-dual-domain decoder
+[…]
 1 Introduction
-In today’s Medical field, drug development is complex and time-consuming.
-Developing a new drug involves three stages: the discovery stage, the preclinical stage, and the clinical stage, which typically takes 10-15 years, with an average cost of up to \$2.8 billion fisher2006alterations ; vijayan2022enhancing .
-Although scientists continue to discover new drug molecules through unremitting efforts and innovation, the traditional drug development model often faces many risks, including the termination of drug development due to drug efficacy, drug safety, or commercial concerns.
-Remarkably, a significant proportion (80-90%) of these projects fail in clinical trials vijayan2022enhancing ; dimasi2010trends .
-Therefore, finding a more efficient and precise method for drug development has become critically important.
-Drug repositioning, a strategy that garnered significant attention, is proposed to explore alternative indications and potential side effects for already-licensed medications dudley2011exploiting ; wang2024landscape .
-Compared with traditional drug research and development methods, this approach offers several advantages, such as shorter time, lower cost, and reduced risk.
-Computational drug repositioning, a favored method among researchers, can significantly reduce the time and energy required to identify candidate drugs.
-Computational drug repositioning methods are divided into four categories cai2023machine ; luo2021biomedical : network diffusion-based methods, machine learning-based methods, matrix factorization-based methods, and deep learning-based methods.
-Network diffusion-based methods classify the drug reposition-ing problem as a link prediction issue within graph neural networks kumar2020link .
-These methods focus on building heterogeneous networks containing multiple types of biological entities and linking drugs to diseases through message propagation on paths bridging different networks.
-A commonly used method is a random walk and its variants luo2021biomedical ; lagcn .
-For instance, Martinez et al.
-martinez2015drugnet proposed a network-based drug repositioning prioritization method called DrugNet.
-It integrates information about diseases, drugs, and proteins to simultaneously prioritize drug-disease and disease-drug associations.
-Another example is Luo et al.’s method BIRW_Cdataset , which effectively predicted potential associations between drugs and diseases by constructing an integrated similarity network between drugs and diseases and combining it with a bidirectional random walk (BiRW).
-Machine learning-based methods classify drug repositioning as a binary classification problem.
-These methods typically use traditional machine learning techniques for drug-disease association prediction, including random forest (RF), support vector machine (SVM), and decision trees cai2023machine .
-For example, Peng et al.
-screening designed a novel DTI screening framework called PUDTI.
-PUDTI utilizes SVM to infer candidate targets for drug repositioning.
+In the field of medicine, drug development is a complex and time-consuming process.
+It involves three main stages: the discovery stage, the preclinical stage, and the clinical stage.
+This process typically takes 10-15 years and around \$2.8 billion on average (Fisher Wilson, 2006; Vijayan et al., 2022).
+Despite continuous efforts and innovation, the traditional drug development model faces significant risks and sometimes leading to the termination of projects due to concerns about drug efficacy, safety, or commercial viability.
+Notably, a high proportion (80-90%) of these projects fail during clinical trials (Vijayan et al., 2022; DiMasi et al., 2010).
+Therefore, finding a more efficient and precise method for drug development is critically important and valuable.
+Computational drug repositioning is a powerful strategy that aims to discover alternative uses of already approved drugs taking into account their potential side effects (Dudley et al., 2011; Wang et al., 2024).
+This approach uses computational and bioinformatics tools and is more efficient and cost-effective than traditional drug research and development methods.
+It allows for rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options.
+Repurposing existing drugs can significantly reduce development costs compared to developing new drugs from scratch.
+In response to new health crises, such as pandemics, repurposing known drugs can provide a quicker response compared to developing new treatments.
+This method can also identify new usage for drugs that may be effective for diseases lacking effective treatments and thereby addressing significant health challenges.
+Machine learning is commonly used in computational drug repositioning.
+Traditional techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose (Cai et al., 2023).
+For instance, Peng et al.
+(2017) introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets.
 Shi et al.
-shi2019lasso proposed a method for predicting drug-target interactions based on Lasso dimensionality reduction and a random forest classifier.
+(2019) proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier.
 Zhao et al.
-HINGRL proposed HINGRL, which uses a random forest classifier in the final decoding step to predict drug-disease association.
-In addition, there are decision trees gradient , linear regression, and logistic regression you2019predicting ; elisei2019recursive .
-Matrix decomposition-based methods categorize the drug repositioning problem as a recommendation system issue.
-These methods employ matrix factorization techniques to reduce dimensionality and extract key features, thus better capturing the associations between drugs and diseases.
-Some standard matrix factorization techniques include singular value decomposition (SVD) and non-negative matrix factorization (NMF).
-For example, Yang et al.
-MSBMF proposed a Multiple Similarity Bilinear Matrix Factorization (MSBMF) method, where the similarity matrices of drugs and diseases are concatenated and the drug-disease association matrix is reshaped through matrix decomposition.
-Luo et al.
-luo2018SVT employed a fast Singular Value Thresholding (SVT) algorithm to propose a drug repositioning recommendation system (DRRS).
-Additionally, Sadegh et al.
-sadeghi2022NMF-DR and Xuan et al.
-xuan2019dd combined non-negative matrix factorization to predict drug-disease associations.
-In recent years, methods based on deep learning have become a hot research topic for drug repositioning.
-They can effectively capture high-level hidden representations from various data sources and demonstrate state-of-the-art performance.
-Unlike traditional machine learning methods, deep learning approaches can learn more abstract information by building deep architectures without manual feature selection and tuning yu2022deep .
-Generally, methods based on deep learning for drug repositioning often integrate graph neural networks such as Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to construct models.
-For example, Yu et al.
-lagcn integrated a drug-disease heterogeneous network with a GCN network to build a multi-layer connected end-to-end network model called LAGCN.
-They also incorporated layer attention mechanisms for drug-disease association prediction.
-Gu et al.
-gu2022REDDA designed a biological network of five biological entities and equipped it with three attention mechanisms to predict drug-disease associations.
-Wang et al.
-wang2023MSGCA designed a multi-similarity graph convolutional autoencoder(MSGCA).
-Sun et al.
-sun2022PSGCN predicted new drug-disease associations by constructing partner-specific graphs for drug-disease pairs and combining them with the GCN.
-By incorporating the GCN, Xuan et al.
-xuan2022heterogeneous constructed multiple drug-disease heterogeneous networks based on different drug similarities and encoded drugs and diseases at Multi-scale neighbors.
+(2022) and Yang et al.
+(2023) employed random forests in their final prediction step, while other techniques such as decision trees (Xuan et al., 2019), matrix factorization(Yang et al., 2021; Luo et al., 2018; Sadeghi et al., 2022), and logistic regression (Yu et al., 2020) have also been developed.
+In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by (Zhang et al., 2017; Mahdaddi et al., 2021; Bansal et al., 2024).
+Unlike traditional machine learning methods, deep learning technology can effectively capture high-level hidden representation using deep architectures without manual feature selection and tuning (Yu et al., 2022).
+Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building (Yu et al., 2021; Cai et al., 2021; Sun et al., 2022; Tang et al., 2023; Zhang et al., 2024).
+For instance, Yu et al.
+(2021) performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases, but did not account for the differences in encoding from different network spaces.
 Cai et al.
-cai2021DRHGCN designed intra-domain and inter-domain feature extraction modules through the GCN and the bilinear aggregator, proposing DRHGCN.
-Meng et al.
-DRAGNN utilized a GAT for encoding drugs and diseases and introduced a weighted local information augmented graph neural network model named DRAGNN.
-Most current deep learning-based methods for drug repositioning typically utilize the similarity between neighbors and aggregate neighbor features for training, achieving considerable results.
-However, many end-to-end deep learning models do not involve purposeful encoding.
-They delegate the task to the optimizer, relying solely on the coordinated training of the encoding and decoding processes to complete association prediction.
-This may lead to the introduction of a large amount of noisy information during model training, resulting in imprecise encoding of drugs and diseases and affecting the prediction results.
-This paper proposes a dual-feature drug repositioning neural network (DFDRNN) model to address the above issue.
-The DFDRNN constructs two features for drugs and diseases: the similarity feature and the association feature.
-It utilizes the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module to extract drug and disease features and achieve precise encoding.
-A self-attention mechanism (SAM) is employed in these two modules to capture the complex relationships between adjacent nodes.
-Finally, the decoder adopts a cross-dual-domain decoder to achieve precise prediction of drug-disease associations.
-Compared to GCN-based methods, DFDRNN employs a SAM to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation.
-This enables efficient aggregation of neighbors and enhances the model’s predictive power.
-We compare DFDRNN with six state-of-the-art methods on four benchmark datasets.
-Experimental results show that our model outperforms other models.
-This article makes the following contributions:
-(1) We propose a dual-feature method to represent drugs and diseases.
-The proposed approach involves two features: the similarity feature and the association feature.
-These two features are independent of each other in the network, while they can be extracted simultaneously to represent drugs and diseases in each network layer.
-(2) Two dual-feature extraction modules are proposed: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
-In particular, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature.
-The proposed two modules effectively capture the features of drugs and diseases, ensuring a precise encoding of drugs and diseases.
-(3) A cross-dual-domain decoder is proposed, implementing a cross-computation between two encoded features to obtain decoding results for the drug and disease domains.
-The fusion of the two decoding results improves the accuracy of association prediction.
+(2021) distinguished between intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases.
+Zhang et al.
+(2024) extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features.
+However, these studies did not consider the calculation of information weights from different neighbors.
+Authors in (Gu et al., 2022; Meng et al., 2024; Huang et al., 2024) used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding.
+Most deep learning models have so far been focused on extracting valuable features from neighborhoods.
+However, these models often employ single encoding approach and lack intentional encoding and decoding processes for association prediction.
+This can lead to the introduction of a large amount of noisy information during model training, resulting in the imprecise encoding of drugs and diseases, ultimately affecting the accuracy of prediction results.
+In order to address this challenge, we propose a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model that uses two features, similarity, and association, to accurately represent drugs and diseases.
+In comparison to existing methods, the proposed model incorporates two dual-feature extraction modules: the IntraDDFE module for extracting features within a single domain (drug or disease) and the InterDDFE module for extracting features across domains.
+DFDRNN uses a self-attention mechanism (SAM) to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation.
+This allows for efficient aggregation of neighbors and enhances the model’s predictive power.
+Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains.
+The key contributions of this paper are:
+(1) To predict potential associations between drugs and diseases, we propose a dual-feature drug repositioning neural network (DFDRNN) model.
+A self-attention mechanism is used capture complex relationship among adjacent nodes.
+The encoder of DFDRNN consists of two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
+(2) The IntraDDFE module and InterDDFE module extract features exploiting information within a single domain and across domains, respectively, and track the dynamic changes of dual-feature to realize precise encoding of drugs and diseases.
+(3) A cross-dual-domain decoder is proposed to perform cross-computation between two encoded features.
+This process yields decoding results for both the drug and disease domains, which help effectively predicting the drug-disease associations.
+(4) The performance of the proposed method is compared with that of six state-of-the-art methods on four benchmark datasets.
+The experimental results show that DFDRNN outperforms other models.
+In addition, case studies on Alzheimer’s and Parkinson’s diseases validated the reliability of DFDRNN in practical applications.
 The remainder of this paper is organized as follows.
-Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results; and Section 4 concludes our work.
+Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results, and Section 4 concludes this work with future research directions.
 2 Materials and Methods
@@
 We conducted experiments on four benchmark datasets.
-The first dataset, Fdataset Fdataset , includes 593 drugs from DrugBank knox2024drugbank and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) hamosh2005OMIM database, comprising 1933 drug-disease associations.
-The second dataset, Cdataset BIRW_Cdataset , consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations.
-The third dataset, Ldataset lagcn , contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) davis2023CTD .
-The fourth dataset, LRSSL liang2017lrssl , comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations.
+The first dataset, Fdataset (Gottlieb et al., 2011), includes 593 drugs from DrugBank (Knox et al., 2024) and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) (Hamosh et al., 2005) database, comprising 1933 drug-disease associations.
+The second dataset, Cdataset (Luo et al., 2016), consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations.
+The third dataset, Ldataset (Yu et al., 2021), contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) (Davis et al., 2023).
+The fourth dataset, LRSSL (Liang et al., 2017), comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations.
 Table 1 summarizes the four datasets’ statistics.
-This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation weininger1988smiles .
-The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit steinbeck2003chemistry .
-Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner van2006text .
-MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database lipscomb2000medical .
+This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation (Weininger, 1988).
+The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit (Steinbeck et al., 2003).
+Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner (Van Driel et al., 2006).
+MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database (Lipscomb, 2000).
 Table 1: The statistics of the four datasets used in this study
-[…]
-[…]
+Datasets No. of drugs No. of disease No. of assocations
+Fdataset 593 313 1933
 Cdataset 663 409 2532
@@
 LRSSL 763 681 3051
-[…]
 2.2 Model structure
@@
 The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
-While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the feature.
-A self-attention (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes.
+While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the features.
+A self-attention mechanism (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes.
 The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains.
@@
 Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks.
-The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on self-attention.
+The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on the self-attention mechanism.
 A layer attention mechanism is also introduced to compute the final similarity and association features.
@@
 2.2.1 Three networks
-To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity of the two drug incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge.
+To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity between the two drugs incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge.
 We use three matrices to represent these three networks.
 Specifically, suppose that there are $n$ drugs (denoted by $r_{1},r_{2},\ldots,r_{n}$) and $m$ diseases (denoted by $d_{1},d_{2},\ldots,d_{m}$).
-We use $S^{r}$ (or $S^{d}$) to denote the adjacency matrix of a drug-drug similarity network (or disease-disease similarity network), which is a $n\times n$ ($m\times m$) real matrix such that the entry $s_{ij}\in[0,1]$ denotes the degree of similarity of drug $i$ and drug $j$ (or disease $i$ and disease $j$), where $i,j\in\{1,2,\ldots,n\}$ (or $i,j\in\{1,2,\ldots,m\}$).
+We use $S^{r}$ (or $S^{d}$) to denote the adjacency matrix of a drug-drug similarity network (or disease-disease similarity network), which is an $n\times n$ ($m\times m$) real matrix such that the entry $s^{r}_{ij}$ (or $s^{d}_{ij}$) $\in[0,1]$ denotes the degree of similarity of drug $i$ and drug $j$ (or disease $i$ and disease $j$), where $i,j\in\{1,2,\ldots,n\}$ (or $i,j\in\{1,2,\ldots,m\}$).
 For a drug-disease association network, we use a $n\times m$ binary matrix, denoted by $A$, to represent it, where the entry $a_{ij}=1$ if drug $i$ is associated with disease $j$ and $a_{ij}=0$ if drug $i$ is not associated with disease $j$ or the associativity is yet to be observed, for $i=1,2,\ldots,n$ and $j=1,2,\ldots,m$.
 To reduce the impact of data noise, we aggregate only the top-$t$ neighbors for each drug in $S^{r}$ and each disease in $S^{d}$, and then two new matrices $R\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{m\times m}$ are generated to replace $S^{r}$ and $S^{d}$, respectively, where the entries $r_{ij}$ ($i,j\in\{1,2,\ldots,n\}$) of $R$ and $d_{ij}$ ($i,j\in\{1,2,\ldots,m\}$) of $D$ are defined as follows.
-For a drug $r_{i}$ (and a disease $d_{i}$), we denote by $N_{k}(r_{i})$ (and $N_{k}(d_{i})$) the set of the $t$-nearest neighbors of $r_{i}$ in $S^{r}$ (and $d_{i}$ in $S^{d}$), i.e., the weight of edge $r_{i}x$ (and edge $d_{i}y$) for every $x\in N_{k}(r_{i})$ (and every $y\in N_{k}(d_{i})$) is top-$t$ among all neighbors of $r_{i}$ (and $d_{i}$), and let $\tilde{N_{k}}(r_{i})=N_{k}(r_{i})\cup\{r_{i}\}$ (and $\tilde{N_{k}}(d_{i})=N_{k}(d_{i})\cup\{d_{i}\}$) be the extended $t$-nearest neighbors of $r_{i}$ (and $d_{i}$).
-$\begin{matrix}R_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{k}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&D_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{k}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)
+For a drug $r_{i}$ (and a disease $d_{i}$), we denote by $N_{t}(r_{i})$ (and $N_{t}(d_{i})$) the set of $t$-nearest neighbors of $r_{i}$ in $S^{r}$ (and $d_{i}$ in $S^{d}$), i.e., the weight of edge $r_{i}x$ (and edge $d_{i}y$) for every $x\in N_{t}(r_{i})$ (and every $y\in N_{t}(d_{i})$) is top-$t$ among all neighbors of $r_{i}$ (and $d_{i}$), and let $\tilde{N_{t}}(r_{i})=N_{t}(r_{i})\cup\{r_{i}\}$ (and $\tilde{N_{t}}(d_{i})=N_{t}(d_{i})\cup\{d_{i}\}$) be the extended $t$-nearest neighbors of $r_{i}$ (and $d_{i}$).
+$\begin{matrix}r_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{t}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&d_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{t}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)
 Finally, to encode drugs and diseases simultaneously and distinguish IntraDDFE from InterDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$.
@@
 2.2.2 Self-attention mechanism
-The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q$ and a group of $K$-$V$ pairs to a new output vector.
-The output of the attention is a weighted sum of $V$ vectors, where the weight for each $V$ is computed by taking the dot product of the $Q$ with the corresponding $K$ vaswani2017attention .
+The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector.
+The output of the attention is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ (Vaswani et al., 2017).
 In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases.
-We use the notation sam to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$.
-The specific implementation of sam can be described as follows:
+We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$.
+The specific implementation of $\mathbf{sam}$ can be described as follows:
 $Q=H^{in}w^{Q},\qquad K=H^{in}w^{K},\qquad V=H^{in}w^{V}$ (3)
@@
 Then, the weight of aggregating neighbor information is calculated based on $Q^{q}$ and $K^{q}$.
-Specifically, for a node $i$ (drug or disease), the weight $\beta^{r}_{ij}$ of $i$ aggregating its neighbor $j$ is defined as
+Specifically, for a node $i$ (drug or disease), the weight $\beta^{q}_{ij}$ of $i$ aggregating its neighbor $j$ is defined as
 $\beta^{q}_{ij}=\frac{exp(\frac{Q^{q}_{i}\bigodot K^{q}_{j}}{\sqrt{k^{\prime}}})}{\sum_{z\in N(i)}exp(\frac{Q^{q}_{i}\bigodot K^{q}_{z})}{\sqrt{k^{\prime}}})}$ (4)
 […]
-$H_{\texttt{sam}}^{i}=\texttt{concat}(\sum_{j\in N(i)}\beta^{1}_{ij}V^{1}_{j},\sum_{j\in N(i)}\beta^{2}_{ij}V^{2}_{j},\ldots,\sum_{j\in N(i)}\beta^{p}_{ij}V^{p}_{j})$ (5)
-Moreover, let $H_{\texttt{sam}}$ represent the feature matrix obtained by SAM, $H_{r}\in\mathbb{R}^{n\times k}$ be the drug feature obtained by SAM, and $H_{d}\in\mathbb{R}^{m\times k}$ is the disease feature obtained by SAM.
+$H_{\mathbf{sam}}^{i}=\mathbf{concat}(\sum_{j\in N(i)}\beta^{1}_{ij}V^{1}_{j},\sum_{j\in N(i)}\beta^{2}_{ij}V^{2}_{j},\ldots,\sum_{j\in N(i)}\beta^{p}_{ij}V^{p}_{j})$ (5)
+Moreover, let $H_{\mathbf{sam}}$ represent the feature matrix obtained by SAM, $H_{r}\in\mathbb{R}^{n\times k}$ be the drug feature obtained by SAM, and $H_{d}\in\mathbb{R}^{m\times k}$ is the disease feature obtained by SAM.
 Then, we have
-$H_{r}=\begin{bmatrix}H^{r_{1}}_{\texttt{sam}}\\ H^{r_{2}}_{\texttt{sam}}\\ \dots\\ H^{r_{m}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\texttt{sam}}\\ H^{d_{2}}_{\texttt{sam}}\\ \dots\\ H^{d_{n}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{\texttt{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
+$H_{r}=\begin{bmatrix}H^{r_{1}}_{\mathbf{sam}}\\ H^{r_{2}}_{\mathbf{sam}}\\ \dots\\ H^{r_{n}}_{\mathbf{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\mathbf{sam}}\\ H^{d_{2}}_{\mathbf{sam}}\\ \dots\\ H^{d_{m}}_{\mathbf{sam}}\end{bmatrix},\hskip 8.5359ptH_{\mathbf{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
 A fully connected layer is applied after the SAM function to extract further features of drugs and diseases.
-We use the notation samf to denote the whole process, i.e., samf is the combination of sam with a fully connected (FC) layer.
-Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU maas2013leakyrelu activation function.
+We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer.
+Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU (Maas et al., 2013) activation function.
 Then,
-$\texttt{samf}(E,H^{in})$ $=\texttt{FC}(\texttt{sam}(E,H^{in}))$ (7)
+$\mathbf{samf}(E,H^{in})$ $=\mathbf{FC}(\mathbf{sam}(E,H^{in}))$ (7)
 $=\begin{bmatrix}\sigma(H_{r}w_{r}+b_{r})\\ \sigma(H_{d}w_{d}+b_{d})\end{bmatrix}$
@@
 Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$.
-In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\texttt{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE.
+In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\mathbf{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE.
 Then,
 $(\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a})$ $=f^{\ell}(E_{s},H^{\ell}_{s},H^{\ell}_{a})$ (10)
-$=(\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
-For the InterDDFE module, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer.
-Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\texttt{samf}^{\ell}_{g}$ represents the aggregation function.
+$=(\mathbf{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\mathbf{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
+Since IntraDDFE extracts features within a single domain, the dual-feature does not undergo dynamic changes.
+Thus, the input similarity features $H^{\ell}_{s}$ produce the output similarity features $\hat{H}^{\ell+1}_{s}$, and the input association features $H^{\ell}_{a}$ produce the output association features $\hat{H}^{\ell+1}_{a}$.
+Next, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer.
+Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\mathbf{samf}^{\ell}_{g}$ represents the aggregation function.
 Then,
 $(\tilde{H}^{\ell+1}_{a},\tilde{H}^{\ell+1}_{s})$ $={g^{\ell}(E_{a},H^{\ell}_{s},H^{\ell}_{a})}$ (11)
-${=(\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
-It should be noted that compared with the IntraDDFE module, the InterDDFE module changes the features: the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$ and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
+${=(\mathbf{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\mathbf{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
+Since the InterDDFE module extracts features across domains, dual-feature undergo dynamic changes.
+Specifically, the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$, and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
 Finally, for each module, the similarity (and association) features (including one input feature and two output features corresponding to the two modules) in the $\ell$-layer are added into an input similarity (and association) feature of the $(\ell+1)$-layer, i.e.,
@@
 That is,
-${s_{r}}$ $=\texttt{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
-${s_{d}}$ $=\texttt{sigmod}(H^{d}_{a}\times(H^{r}_{s})^{T})$ (16)
+${s_{r}}$ $=\mathbf{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
+${s_{d}}$ $=\mathbf{sigmod}(H^{d}_{a}\times(H^{r}_{s})^{T})$ (16)
 ${s}$ $={(s_{r}+(s_{d})^{T})/2.0}$ (17)
@@
 3:  Initialize the weight matrix ($w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$).
-4:for $c=1$ to $C$ do
+4:   for $c=1$ to $C$
 5:        Compute the similarity feature $H^{0}_{s}$ and the association feature $H^{0}_{a}$ with Equation (9).
-6:     for $i=1$ to $L$ do
+6:        for $i=1$ to $L$
 7:              Compute IntraDDFE $\hat{H}^{\ell}_{s},\hat{H}^{\ell}_{a}$ with Equation (10).
@@
 We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively.
-We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($i$)-disease($j$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
-$loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{i,j})+\sum_{(i,j)\in y^{-}}\log(1-s_{i,j}))$ (18)
-We optimize the DFDRNN model using the Adam kingma2014adam optimizer and initialize the parameters’ weights with Xavier initialization glorot2010understanding .
-We randomly drop some edges and apply regular dropout srivastava2014dropout to improve the model’s generalization ability.
+We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($r_{i}$)-disease($d_{j}$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
+$loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{ij})+\sum_{(i,j)\in y^{-}}\log(1-s_{ij}))$ (18)
+We optimize the DFDRNN model using the Adam (Kingma and Ba, 2014) optimizer and initialize the parameters’ weights with Xavier initialization (Glorot and Bengio, 2010).
+We randomly drop some edges and apply regular dropout (Srivastava et al., 2014) to improve the model’s generalization ability.
 3 Experiments and Results
@@
 To assess DFDRNN’s performance, we compared it with six state-of-the-art drug-disease association prediction models across the above four datasets. These models are:
-$\bullet$ LAGCN lagcn : This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
-$\bullet$ DRHGCN cai2021DRHGCN : This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
-$\bullet$ DRWBNCF meng2022DRWBNCF : This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
-$\bullet$ NCH-DDA zhang2024NCH-DDA : This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases.
+$\bullet$ LAGCN (Yu et al., 2021): This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
+$\bullet$ DRHGCN (Cai et al., 2021): This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
+$\bullet$ DRWBNCF (Meng et al., 2022): This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
+$\bullet$ NCH-DDA (Zhang et al., 2024): This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases.
 Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
-$\bullet$ HDGAT huang2024HDGAT : This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
-$\bullet$ DRGBCN tang2023DRGBCN : This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
+$\bullet$ HDGAT (Huang et al., 2024): This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
+$\bullet$ DRGBCN (Tang et al., 2023): This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
 3.2 Experimental parameter configuration
@@
 3.3 Evaluation metrics
-Table 2: The statistics of the four datasets used in this study
+We categorize the prediction results into four sets.
+As shown in Table 2, TP represents the number of true positives correctly classified as positive, FN represents the number of false negatives incorrectly classified as negative, FP represents the number of false positives incorrectly classified as positive, and TN represents the number of true negatives correctly classified as negative.
+We then calculate True Positive Rate (TPR or Recall), False Positive Rate (FPR), and precision as follows:
+Table 2: The statistics of the four sets used in this study
 Actual Predicted Positive Negative
@@
 Negative FP TN
-We categorize the prediction results into four sets.
-As shown in Table 2, TP represents the number of true positives correctly classified as positive, FN represents the number of false negatives incorrectly classified as negative, FP represents the number of false positives incorrectly classified as positive, and TN represents the number of true negatives correctly classified as negative.
-We then calculate True Positive Rate (TPR or Recall), False Positive Rate (FPR), and precision as follows:
 […]
@@
 Table 3: The AUROCs and AUPRs of DFDRNN and the other comparative models under 10-times 10-fold cross-validation.
-[…]
+Metrics Model Fdataset Cdataset Ldataset LRSSL Avg.
 AUROC LAGCN 0.881±0.002 0.911±0.002 0.864±0.001 0.935±0.001 0.898
@@
 DFDRNN 0.623±0.003 0.688±0.002 0.573±0.001 0.502±0.002 0.597
-[…]
-{tablenotes}
-The best results in each row are in bold faces and the second best results are underlined.
+  • The best results in each column are in bold faces and the second best results are underlined.
 3.5 Effect of the number of nearest neighbors
@@
 Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$.
-However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention aggregation capability.
-Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across four datasets.
+However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention mechanism’s aggregation capability.
+Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across the four datasets.
 For detailed numerical values, please refer to Supplementary Tables S1 and S2.
@@
   • DFDRNN-noCF: Remove the change of features in InterDDFE, meaning that the similarity feature and the association feature will not be changed into each other after InterDDFE.
+Table 4: The AUROCs and AUPRs of DFDRNN and its four variants under 10-fold cross-validation
+Variants Fdataset Cdataset Ldataset LRSSL
+AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
+DFDRNN-SF 0.958 0.612 0.972 0.690 0.873 0.538 0.964 0.464
+DFDRNN-AF 0.942 0.606 0.962 0.678 0.884 0.570 0.950 0.504
+DFDRNN-GCN 0.950 0.620 0.964 0.671 0.874 0.551 0.956 0.483
+DFDRNN-noCF 0.959 0.611 0.970 0.676 0.868 0.523 0.963 0.462
+DFDRNN 0.960 0.622 0.973 0.689 0.885 0.572 0.965 0.504
 Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-fold cross-validation.
@@
 Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Fdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the InterDDFE module.
-Table 4: The AUROCs and AUPRs of DFDRNN and its four variants under 10-fold cross-validation
-[…]
-AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
-DFDRNN-SF 0.958 0.612 0.972 0.690 0.873 0.538 0.964 0.464
-DFDRNN-AF 0.942 0.606 0.962 0.678 0.884 0.570 0.950 0.504
-DFDRNN-GCN 0.950 0.620 0.964 0.671 0.874 0.551 0.956 0.483
-DFDRNN-noCF 0.959 0.611 0.970 0.676 0.868 0.523 0.963 0.462
-DFDRNN 0.960 0.622 0.973 0.689 0.885 0.572 0.965 0.504
-[…]
 3.6.2 Decoder
@@
 The specific implementation of the non-cross dual-domain decoder is described as follows.
-${s^{non}_{r}}$ $=\texttt{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
-${s^{non}_{d}}$ $=\texttt{sigmod}(H^{d}_{s}\times(H^{r}_{s})^{T})$ (23)
+${s^{non}_{r}}$ $=\mathbf{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
+${s^{non}_{d}}$ $=\mathbf{sigmod}(H^{d}_{s}\times(H^{r}_{s})^{T})$ (23)
 ${s^{non}}$ $={(s^{non}_{r}+(^{non}_{d})^{T})/2.0}$ (24)
@@
 In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment.
-Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic singh2020drug .
+Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic (Singh et al., 2020).
 Finding suitable drugs for emerging diseases is a complex and crucial task.
@@
 In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated.
-The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCHDDA by 0.99%.
+The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCH-DDA by 0.99%.
 Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
@@
 Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD.
-To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB knox2024drugbank , PubChem kim2016pubchem , CTD davis2023CTD , DrugCentral avram2021drugcentral , and ClinicalTrials.
+To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB (Knox et al., 2024), PubChem (Kim et al., 2016), CTD (Davis et al., 2023), DrugCentral (Avram et al., 2021), and ClinicalTrials.
 AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder.
@@
 It can also be utilized for the prevention or symptomatic treatment of influenza A.
-Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment erkulwater1989amantadine .
+Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment (Erkulwater and Pillai, 1989).
 Carbamazepine functions as an anticonvulsant and analgesic drug.
-Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients lemke1995effect .
+Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients (Lemke, 1995).
 DFDRNN predicts an association with carbamazepine, which has been validated by CTD.
@@
 Table 5: The top 10 DFDRNN-predicted candidate drugs for AD
-[…]
-[…]
+Rank Drug ID Drug Name Evidence
+1 DB00915 Amantadine DB/CTD/DrugCentral/ PubChem/ClinicalTrials
 2 DB00564 Carbamazepine CTD
@@
 10 DB01262 Decitabine CTD
-[…]
 Table 6: The top 10 DFDRNN-predicted candidate drugs for PD
-[…]
-[…]
+Rank Drug ID Drug Name Evidence
+1 DB00502 Haloperidol DB/CTD/PubChem/ ClinicalTrials
 2 DB01068 Clonazepam CTD/ClinicalTrials
-3 DB01235 Levodopa DB/CTD/PubChem/ DrugCentral/ClinicalTrials/
-4 DB00810 Biperiden DB/CTD/PubChem/ DrugCentral
+3 DB01235 Levodopa DB/CTD/DrugCentral/ PubChem/ClinicalTrials
+4 DB00810 Biperiden DB/CTD/DrugCentral/ PubChem
 5 DB00424 Hyoscyamine DB
 6 DB00575 Clonidine ClinicalTrials
-7 DB00915 Amantadine DB/PubChem/DrugCentral/ ClinicalTrials
-8 DB00989 Rivastigmine DB/CTD/PubChem/ DrugCentral/ClinicalTrials
-9 DB00376 Trihexyphenidyl DB/CTD/PubChem/ DrugCentral
+7 DB00915 Amantadine DB/DrugCentral/PubCh-em/ClinicalTrials
+8 DB00989 Rivastigmine DB/CTD/DrugCentral/ PubChem/ClinicalTrials
+9 DB00376 Trihexyphenidyl DB/CTD/DrugCentral/ PubChem
 10 DB00215 Citalopram CTD
-[…]
 Parkinson’s disease (PD) is the second most common neurogenic disorder after AD.
@@
 Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources.
-The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provide reassurance about the reliability of the model.
+The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
 4 Conclusion
@@
 The experimental results confirm the effectiveness of our proposed DFDRNN model in drug repositioning.
-In the future, we will consider incorporating more biological entities, such as miRNA and proteins, for drug repositioning.
-Additionally, we will explore more accurate methods to calculate drug (or disease) similarity to reduce data noise.
-Key Points We propose DFDRNN, a new deep learning method for drug repositioning, which simultaneously extracts two kinds of features: the similarity feature and the association feature to represent drugs and diseases in each network layer.
-Based on two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module, DFDRNN captures the potential relationships between the features of drugs and diseases.
-A cross-dual-domain decoder is proposed to implement a cross-computation between two encoded features to redict drug-disease associations.
-DFDRNN outperformed the state-of-the-art methods on four benchmark datasets.
-5 Acknowledgments
+For future work, we would like to investigate whether specific enhancements, such as integrating temporal data and handling multi-modal data, could further enhance the performance of DFDRNN.
+By incorporating temporal data, the model can better understand and predict patterns over time.
+Integration of data from different modalities could provide a more comprehensive understanding of the input.
+We aim to identify and implement the appropriate techniques and models for these enhancements.
+Furthermore, we intend to explore the integration of additional biological entities, such as miRNA and proteins, for drug repositioning.
+Acknowledgments
 This work was supported by the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China under grants 62172072 and 62272115.

### v3

@@
-Conceptualization of this study, Methodology, Software
+Boosting drug-disease association prediction for drug repositioning via dual-feature extraction and cross-dual-domain decoding
 Abstract
-Uncovering new therapeutic uses of existing drugs, drug repositioning offers a fast and cost-effective strategy and holds considerable significance in the realm of drug discovery and development.
-In recent years, deep learning techniques have emerged as powerful tools in drug repositioning due to their ability to analyze large and complex datasets.
-However, many existing methods focus on extracting feature information from nearby nodes in the network to represent drugs and diseases, without considering the potential inter-relationships between the features of drugs and diseases, which may lead to inaccurate representations.
-To address this limitation, we use two features (similarity and association) to capture the potential relationships between the features of drugs and diseases, proposing a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model.
-DFDRNN uses a self-attention mechanism to extract neighbor features and incorporates two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module for extracting features within a single domain (drugs or diseases) and the inter-domain dual-feature extraction (InterDDFE) module for extracting features across domains.
+The extraction of biomedical data has significant academic and practical value in contemporary biomedical sciences.
+In recent years, drug repositioning, a cost-effective strategy for drug development by discovering new indications for approved drugs, has gained increasing attention.
+However, many existing drug repositioning methods focus on mining information from adjacent nodes in biomedical networks without considering the potential inter-relationships between the feature spaces of drugs and diseases.
+This can lead to inaccurate encoding, resulting in biased mined drug-disease association information.
+To address this limitation, we propose a new model called Dual-Feature Drug Repurposing Neural Network (DFDRNN).
+DFDRNN allows the mining of two features (similarity and association) from the drug-disease biomedical networks to encode drugs and diseases.
+A self-attention mechanism is utilized to extract neighbor feature information.
+It incorporates two dual-feature extraction modules: the single-domain dual-feature extraction (SDDFE) module for extracting features within a single domain (drugs or diseases) and the cross-domain dual-feature extraction (CDDFE) module for extracting features across domains.
 By utilizing these modules, we ensure more appropriate encoding of drugs and diseases.
-Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains.
+A cross-dual-domain decoder is also designed to predict drug-disease associations in both domains.
 Our proposed DFDRNN model outperforms six state-of-the-art methods on four benchmark datasets, achieving an average AUROC of 0.946 and an average AUPR of 0.597.
 Case studies on two diseases show that the proposed DFDRNN model can be applied in real-world scenarios, demonstrating its significant potential in drug repositioning.
-keywords:
-[…]
-1 Introduction
-In the field of medicine, drug development is a complex and time-consuming process.
-It involves three main stages: the discovery stage, the preclinical stage, and the clinical stage.
-This process typically takes 10-15 years and around \$2.8 billion on average (Fisher Wilson, 2006; Vijayan et al., 2022).
-Despite continuous efforts and innovation, the traditional drug development model faces significant risks and sometimes leading to the termination of projects due to concerns about drug efficacy, safety, or commercial viability.
-Notably, a high proportion (80-90%) of these projects fail during clinical trials (Vijayan et al., 2022; DiMasi et al., 2010).
-Therefore, finding a more efficient and precise method for drug development is critically important and valuable.
-Computational drug repositioning is a powerful strategy that aims to discover alternative uses of already approved drugs taking into account their potential side effects (Dudley et al., 2011; Wang et al., 2024).
-This approach uses computational and bioinformatics tools and is more efficient and cost-effective than traditional drug research and development methods.
-It allows for rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options.
-Repurposing existing drugs can significantly reduce development costs compared to developing new drugs from scratch.
-In response to new health crises, such as pandemics, repurposing known drugs can provide a quicker response compared to developing new treatments.
-This method can also identify new usage for drugs that may be effective for diseases lacking effective treatments and thereby addressing significant health challenges.
-Machine learning is commonly used in computational drug repositioning.
-Traditional techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose (Cai et al., 2023).
-For instance, Peng et al.
-(2017) introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets.
-Shi et al.
-(2019) proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier.
-Zhao et al.
-(2022) and Yang et al.
-(2023) employed random forests in their final prediction step, while other techniques such as decision trees (Xuan et al., 2019), matrix factorization(Yang et al., 2021; Luo et al., 2018; Sadeghi et al., 2022), and logistic regression (Yu et al., 2020) have also been developed.
-In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by (Zhang et al., 2017; Mahdaddi et al., 2021; Bansal et al., 2024).
-Unlike traditional machine learning methods, deep learning technology can effectively capture high-level hidden representation using deep architectures without manual feature selection and tuning (Yu et al., 2022).
-Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building (Yu et al., 2021; Cai et al., 2021; Sun et al., 2022; Tang et al., 2023; Zhang et al., 2024).
-For instance, Yu et al.
-(2021) performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases, but did not account for the differences in encoding from different network spaces.
-Cai et al.
-(2021) distinguished between intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases.
-Zhang et al.
-(2024) extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features.
-However, these studies did not consider the calculation of information weights from different neighbors.
-Authors in (Gu et al., 2022; Meng et al., 2024; Huang et al., 2024) used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding.
-Most deep learning models have so far been focused on extracting valuable features from neighborhoods.
-However, these models often employ single encoding approach and lack intentional encoding and decoding processes for association prediction.
-This can lead to the introduction of a large amount of noisy information during model training, resulting in the imprecise encoding of drugs and diseases, ultimately affecting the accuracy of prediction results.
-In order to address this challenge, we propose a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model that uses two features, similarity, and association, to accurately represent drugs and diseases.
-In comparison to existing methods, the proposed model incorporates two dual-feature extraction modules: the IntraDDFE module for extracting features within a single domain (drug or disease) and the InterDDFE module for extracting features across domains.
+Index Terms:
+Drug Repositioning, Self-Attention Mechanism, Dual-Feature Extraction, Cross-Dual-Domain
+I Introduction
+Drug development is a significant yet complex and time-consuming process.
+It consists of three main stages: the discovery stage, the preclinical stage, and the clinical stage.
+Typically, the entire process takes about 10 to 15 years and costs around \$2.8 billion on average [1, 2].
+Despite ongoing efforts and innovations, the traditional drug development model faces considerable risks, often resulting in project terminations due to concerns about drug efficacy, safety, or commercial viability.
+Notably, a high percentage (80-90%) of these projects fail during clinical trials.
+Consequently, finding a more efficient and precise method for drug development is critically important and valuable.
+Given the high cost of drug development, computational drug repositioning has become a powerful strategy for discovering alternative uses of already approved drugs while considering their potential effects [3].
+This approach utilizes computational and bioinformatics tools, providing greater efficiency and cost-effectiveness compared to traditional drug research and development methods.
+It enables the rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options.
+Repurposing existing drugs can significantly lower development costs when compared to creating new drugs from scratch.
+In response to health crises, such as pandemics, repurposing known drugs offers a quicker solution than developing entirely new treatments.
+Furthermore, this method can unveil new applications for drugs that may be effective against diseases that currently lack effective therapies, thereby addressing major health challenges.
+Machine learning has been commonly used in computational drug repositioning by information extraction from biomedical data.
+Techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose [4].
+For instance, [5] introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets.
+[6] proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier.
+[7] employed random forests in their final prediction step, while other techniques such as decision trees [8], matrix factorization[9, 10], and logistic regression [11] have also been developed.
+In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by [12, 13, 14, 15].
+Unlike traditional machine learning methods, deep learning techniques leverage deep architectures to effectively capture high-level latent representations without requiring manual feature selection and tuning [16].
+Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building [17, 18, 19, 20, 21].
+For instance, [17] performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases but did not account for the differences in encoding from different network spaces.
+[18] distinguished intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases.
+[21] extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features.
+However, these studies did not consider calculating information weights from different neighbors.
+Authors in [22, 23] used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding.
+Most deep learning methods have primarily focused on extracting valuable features from neighborhoods.
+However, these approaches often rely on a single encoding method, which does not intentionally account for relation prediction.
+As a result, a significant amount of noise can arise during model training, hindering the effective extraction of drug-disease association information and ultimately impacting the accuracy of the prediction results.
+To address this challenge, we propose a new model called Dual-Feature Drug Repositioning Neural Network (DFDRNN).
+DFDRNN simultaneously uses two features, similarity and association, to encode drugs and diseases accurately.
+Compared to existing methods, the proposed model incorporates two dual-feature extraction modules: the SDDFE module for extracting features within a single domain (drug or disease) and the CDDFE module for extracting features across domains.
 DFDRNN uses a self-attention mechanism (SAM) to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation.
 This allows for efficient aggregation of neighbors and enhances the model’s predictive power.
-Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains.
-The key contributions of this paper are:
+A cross-dual-domain decoder is also designed to predict drug-disease associations in both domains.
+The key contributions of this work are:
 (1) To predict potential associations between drugs and diseases, we propose a dual-feature drug repositioning neural network (DFDRNN) model.
-A self-attention mechanism is used capture complex relationship among adjacent nodes.
-The encoder of DFDRNN consists of two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
-(2) The IntraDDFE module and InterDDFE module extract features exploiting information within a single domain and across domains, respectively, and track the dynamic changes of dual-feature to realize precise encoding of drugs and diseases.
+A self-attention mechanism is used to capture complex relationships among adjacent nodes.
+The encoder of DFDRNN consists of two dual-feature extraction modules: the single-domain dual-feature extraction (SDDFE) module and the cross-domain dual-feature extraction (CDDFE) module.
+(2) The SDDFE module and the CDDFE module extract information from similarity networks and association networks, respectively, and track the dynamic changes of dual-feature to achieve precise encoding of drugs and diseases.
 (3) A cross-dual-domain decoder is proposed to perform cross-computation between two encoded features.
-This process yields decoding results for both the drug and disease domains, which help effectively predicting the drug-disease associations.
+This process yields decoding results for both the drug and disease domains, which help effectively predict the drug-disease associations.
 (4) The performance of the proposed method is compared with that of six state-of-the-art methods on four benchmark datasets.
-The experimental results show that DFDRNN outperforms other models.
+The experimental results show that DFDRNN outperforms other methods.
 In addition, case studies on Alzheimer’s and Parkinson’s diseases validated the reliability of DFDRNN in practical applications.
-The remainder of this paper is organized as follows.
-Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results, and Section 4 concludes this work with future research directions.
-2 Materials and Methods
-2.1 Datasets
-We conducted experiments on four benchmark datasets.
-The first dataset, Fdataset (Gottlieb et al., 2011), includes 593 drugs from DrugBank (Knox et al., 2024) and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) (Hamosh et al., 2005) database, comprising 1933 drug-disease associations.
-The second dataset, Cdataset (Luo et al., 2016), consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations.
-The third dataset, Ldataset (Yu et al., 2021), contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) (Davis et al., 2023).
-The fourth dataset, LRSSL (Liang et al., 2017), comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations.
-Table 1 summarizes the four datasets’ statistics.
-This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation (Weininger, 1988).
-The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit (Steinbeck et al., 2003).
-Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner (Van Driel et al., 2006).
-MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database (Lipscomb, 2000).
-Table 1: The statistics of the four datasets used in this study
+The rest of this paper is structured as follows: Section II introduces the datasets and methods, providing details on the DFDRNN implementation; Section III explains the experimental setup and discusses the results; Section IV concludes the study and suggests future research directions.
+II Materials and Methods
+II-A Datasets
+We performed experiments on four benchmark datasets.
+The first, Gdataset [24], comprises 593 drugs from DrugBank [25] and 313 diseases from the Online Mendelian Inheritance in Man (OMIM) [26], with a total of 1933 drug-disease associations.
+The second dataset, Cdataset [27], consists of 663 drugs listed in DrugBank, 409 diseases from OMIM, and 2532 known associations.
+Ldataset [17], the third dataset, contains 18,416 associations between 269 drugs and 598 diseases, derived from the Comparative Toxicogenomics Database (CTD) [28].
+Lastly, LRSSL [29] contains 763 drugs, 681 diseases, and 3051 validated drug-disease associations.
+A summary of the statistics for these datasets can be found in Table I.
+This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation [30].
+For disease similarity, it is determined based on disease phenotypes and computed via MimMiner [31], which measures similarity by analyzing the MeSH terms in the medical descriptions of diseases found in the OMIM database [32].
+TABLE I: Summary statistics of the four datasets
 Datasets No. of drugs No. of disease No. of assocations
-Fdataset 593 313 1933
+Gdataset 593 313 1933
 Cdataset 663 409 2532
@@
 LRSSL 763 681 3051
-2.2 Model structure
+II-B Model structure
 Our research is focused on predicting potential associations between drugs and diseases.
 To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model.
-The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
-While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the features.
+The encoder of DFDRNN encompasses two dual-feature extraction modules: the single domain dual-feature extraction (SDDFE) module and the cross domain dual-feature extraction (CDDFE) module.
+While the SDDFE module extracts features among drugs (or diseases), the CDDFE module extracts features between drugs and diseases and changes the features.
 A self-attention mechanism (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes.
@@
 Figure 1 illustrates the diagram of the proposed DFDRNN model.
-Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks.
-The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on the self-attention mechanism.
-A layer attention mechanism is also introduced to compute the final similarity and association features.
-Finally, the decoder calculates drug-disease association scores by decoding separately for the drug and disease domains.
-2.2.1 Three networks
+Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases.
+The encoding process employs a multi-layer connected network architecture, where each layer performs SDDFE and CDDFE, while tracking the dynamic changes of dual-feature.
+A layer attention mechanism computes final features.
+The decoder then calculates drug-disease association scores separately for each domain.
+II-B1 Three key networks
 To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity between the two drugs incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge.
@@
 $\begin{matrix}r_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{t}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&d_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{t}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)
-Finally, to encode drugs and diseases simultaneously and distinguish IntraDDFE from InterDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$.
-$E_{s}$ is used for IntraDDFE, while $E_{a}$ is used for InterDDFE, defined as:
+Finally, to encode drugs and diseases simultaneously and distinguish SDDFE from CDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$.
+$E_{s}$ is used for SDDFE, while $E_{a}$ is used for CDDFE, defined as:
 $\ \qquad\qquad E_{s}=\begin{bmatrix}R&0\\ 0&D\end{bmatrix}\qquad E_{a}=\begin{bmatrix}0&A\\ A^{T}&0\end{bmatrix}$ (2)
-2.2.2 Self-attention mechanism
+II-B2 Self-attention mechanism
 The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector.
-The output of the attention is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ (Vaswani et al., 2017).
-In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases.
+The output of the SAM is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ [33].
+In our SDDFE and CDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases.
 We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$.
@@
 $H_{r}=\begin{bmatrix}H^{r_{1}}_{\mathbf{sam}}\\ H^{r_{2}}_{\mathbf{sam}}\\ \dots\\ H^{r_{n}}_{\mathbf{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\mathbf{sam}}\\ H^{d_{2}}_{\mathbf{sam}}\\ \dots\\ H^{d_{m}}_{\mathbf{sam}}\end{bmatrix},\hskip 8.5359ptH_{\mathbf{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
-A fully connected layer is applied after the SAM function to extract further features of drugs and diseases.
+After the SAM function, a fully connected layer is employed to further extract the features of drugs and diseases.
 We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer.
-Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU (Maas et al., 2013) activation function.
+Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU [34] activation function.
 Then,
@@
 $=\begin{bmatrix}\sigma(H_{r}w_{r}+b_{r})\\ \sigma(H_{d}w_{d}+b_{d})\end{bmatrix}$
-2.2.3 Encoder
-The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: IntraDDFE module based on network $E_{s}$ and InterDDFE module based on network $E_{a}$.
+II-B3 Encoder
+The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: SDDFE module based on network $E_{s}$ and CDDFE module based on network $E_{a}$.
 First, the initial features of drugs and diseases are generated in two features, including the similarity feature $H^{init}_{s}$ and the association feature $H^{init}_{a}$, described as follows.
@@
 Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$.
-In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\mathbf{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE.
+In the $\ell$-th layer, we denote by $f^{\ell}$ the SDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by SDDFE, and $\mathbf{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for SDDFE.
 Then,
@@
 $=(\mathbf{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\mathbf{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
-Since IntraDDFE extracts features within a single domain, the dual-feature does not undergo dynamic changes.
+Since SDDFE extracts features within a single domain, the dual-feature does not undergo dynamic changes.
 Thus, the input similarity features $H^{\ell}_{s}$ produce the output similarity features $\hat{H}^{\ell+1}_{s}$, and the input association features $H^{\ell}_{a}$ produce the output association features $\hat{H}^{\ell+1}_{a}$.
-Next, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer.
-Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\mathbf{samf}^{\ell}_{g}$ represents the aggregation function.
+Next, we use $g^{\ell}$ to denote the CDDFE function in the $\ell$-th layer.
+Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by CDDFE, respectively, and $\mathbf{samf}^{\ell}_{g}$ represents the aggregation function.
 Then,
@@
 ${=(\mathbf{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\mathbf{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
-Since the InterDDFE module extracts features across domains, dual-feature undergo dynamic changes.
+Since the CDDFE module extracts features across domains, dual-feature undergo dynamic changes.
 Specifically, the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$, and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
@@
 As the depth of network layers increases, the valuable information for predicting drug-disease association extracted by the DFDRNN model decreases.
-Therefore, we introduce a layer attention mechanism to allocate different weights to the results of different layers.
+Consequently, a layer attention mechanism is incorporated to assign distinct weights to the outputs of each layer.
 We use $L$ to represent the total number of layers in DFDRNN.
@@
 $\begin{bmatrix}H^{r}_{s}\\ H^{d}_{s}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{s}),\qquad\begin{bmatrix}H^{r}_{a}\\ H^{d}_{a}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{a})$ (14)
-2.2.4 Decoder
+II-B4 Decoder
 A cross-dual-domain decoder is utilized to predict drug-disease association.
@@
 Let $s\in\mathbb{R}^{n\times m}$ denote the final score matrix of drug-disease associations, where $s_{ij}$ represents the prediction score between drug $r_{i}$ and disease $d_{j}$.
-That is,
+That is
 ${s_{r}}$ $=\mathbf{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
@@
 5:        Compute the similarity feature $H^{0}_{s}$ and the association feature $H^{0}_{a}$ with Equation (9).
-6:        for $i=1$ to $L$
-7:              Compute IntraDDFE $\hat{H}^{\ell}_{s},\hat{H}^{\ell}_{a}$ with Equation (10).
-8:              Compute InterDDFE $\tilde{H}^{\ell}_{s},\tilde{H}^{\ell}_{a}$ with Equation (11).
-9:              Fuse the same feature and the output of the previous layer yields $H^{\ell}_{s}$ and $H^{\ell}_{a}$ with Equation (12) (13)
+6:        for $\ell=$ to $L-1$
+7:              Compute SDDFE $\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a}$ with Equation (10).
+8:              Compute CDDFE $\tilde{H}^{\ell+1}_{s},\tilde{H}^{\ell+1}_{a}$ with Equation (11).
+9:              Fuse the same feature and the input feature yields $H^{\ell+1}_{s}$ and $H^{\ell+1}_{a}$ with Equation (12) (13)
 10:        end for
@@
 15:   end for
-2.3 Optimization
-Predicting the drug-disease association can be viewed as a binary classification task.
-The objective is to determine whether a given drug is associated with a specific disease.
-Given the serious imbalance between the number of positive samples (drug-disease associations) and negative samples (drug-disease non-associations), we employ binary cross-entropy as the loss function.
-We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively.
-We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($r_{i}$)-disease($d_{j}$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
+II-C Optimization
+Predicting drug-disease associations can be treated as a binary classification problem, where the goal is to determine if a specific drug is associated with a specific disease.
+Due to the significant imbalance between the number of positive samples (associations) and negative samples (non-associations), we utilize binary cross-entropy as the loss function.
+Let $y^{+}$ and $y^{-}$ represent the positive and negative labels in the dataset, respectively.
+To mitigate the effects of this imbalance, we introduce a weight $\lambda=\frac{|y^{-}|}{|y^{+}|}$, where $|y^{-}|$ and $|y^{+}|$ denote the sizes of the negative and positive sample sets.
+A drug-disease pair $(i,j)$ corresponds to $r_{i}$ (drug) and $d_{j}$ (disease) in either $y^{+}$ or $y^{-}$, respectively.
 $loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{ij})+\sum_{(i,j)\in y^{-}}\log(1-s_{ij}))$ (18)
-We optimize the DFDRNN model using the Adam (Kingma and Ba, 2014) optimizer and initialize the parameters’ weights with Xavier initialization (Glorot and Bengio, 2010).
-We randomly drop some edges and apply regular dropout (Srivastava et al., 2014) to improve the model’s generalization ability.
-3 Experiments and Results
-3.1 Baseline methods
-To assess DFDRNN’s performance, we compared it with six state-of-the-art drug-disease association prediction models across the above four datasets. These models are:
-$\bullet$ LAGCN (Yu et al., 2021): This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
-$\bullet$ DRHGCN (Cai et al., 2021): This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
-$\bullet$ DRWBNCF (Meng et al., 2022): This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
-$\bullet$ NCH-DDA (Zhang et al., 2024): This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases.
+The DFDRNN model is optimized using the Adam optimizer [35], with parameters initialized through Xavier initialization [36].
+To enhance generalization, we randomly drop edges and apply regular dropout techniques [37].
+III Experiments and Results
+III-A Baseline methods
+We compared DFDRNN against six state-of-the-art methods for predicting drug-disease associations across the above four datasets. These methods include:
+$\bullet$ LAGCN [17]: This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
+$\bullet$ DRHGCN [18]: This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
+$\bullet$ DRWBNCF [38]: This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
+$\bullet$ NCH-DDA [21]: This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases.
 Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
-$\bullet$ HDGAT (Huang et al., 2024): This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
-$\bullet$ DRGBCN (Tang et al., 2023): This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
-3.2 Experimental parameter configuration
+$\bullet$ HDGAT [23]: This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
+$\bullet$ DRGBCN [20]: This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
+III-B Experimental parameter configuration
 The model’s parameter values are as follows: DFDRNN’s feature embedding dimension is 128; the number of network layers is 3; the number of attention heads is p = 2; the dropout rate is 0.4; the edge dropout rate is 0.5 on the Ldataset and 0.2 on other datasets; the learning rate is 0.008; and the maximum training epoch is 800.
 Additionally, we set top-$t$ to 7 based on parameter sensitivity experiments.
-3.3 Evaluation metrics
-We categorize the prediction results into four sets.
-As shown in Table 2, TP represents the number of true positives correctly classified as positive, FN represents the number of false negatives incorrectly classified as negative, FP represents the number of false positives incorrectly classified as positive, and TN represents the number of true negatives correctly classified as negative.
-We then calculate True Positive Rate (TPR or Recall), False Positive Rate (FPR), and precision as follows:
-Table 2: The statistics of the four sets used in this study
-Actual Predicted Positive Negative
-Positive TP FN
-Negative FP TN
-[…]
-$\text{FPR}=\frac{FP}{FP+TN}$ (20)
-$\text{Precision}=\frac{TP}{TP+FP}$ (21)
-Calculating the FPR, TPR, and precision at various thresholds, we construct the receiver operating characteristic (ROC) curve, with FPR on the X-axis and TPR on the Y-axis.
-Additionally, we generate the precision-recall (PR) curve, with recall on the X-axis and precision on the Y-axis.
-The evaluation metrics used for DFDRNN are the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
-3.4 Performance of DFDRNN in the cross-validation
-We employ the 10-fold cross-validation method to assess the model’s performance.
-During this process, the dataset is randomly partitioned into ten mutually exclusive subsets, each containing 10% known drug-disease association pairs and 10% unknown pairs.
-One subset serves as the test set, while the remaining nine subsets are used to train the model.
-This process is repeated ten times (i.e., 10-fold cross-validation) with different random seeds, and the average and standard deviation of the results are calculated to compare the model’s performance.
-The experimental results, including AUROC and AUPR values, are presented in Table 3.
-The ROC and PR curves can also be found in Supplementary Figures S1-S4.
-In Table 3, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets.
-This performance surpassed the second-ranked NCH-DDA model by 0.75% and 3.1% in AUROC and AUPR, respectively.
-DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations.
-It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
-Table 3: The AUROCs and AUPRs of DFDRNN and the other comparative models under 10-times 10-fold cross-validation.
-Metrics Model Fdataset Cdataset Ldataset LRSSL Avg.
+TABLE II: AUROC and AUPR results for DFDRNN and comparative models using 10-times repeated 10-fold cross-validation
+Metrics Model Gdataset Cdataset Ldataset LRSSL Avg.
 AUROC LAGCN 0.881±0.002 0.911±0.002 0.864±0.001 0.935±0.001 0.898
@@
 DFDRNN 0.623±0.003 0.688±0.002 0.573±0.001 0.502±0.002 0.597
-  • The best results in each column are in bold faces and the second best results are underlined.
-3.5 Effect of the number of nearest neighbors
+  • The highest values in each column are highlighted in bold, while the second-highest values are underlined
+III-C Evaluation metrics
+We classify the prediction outcomes into four categories.
+As illustrated in Table III, TP denotes true positives that are correctly identified as positive, FN represents false negatives that are wrongly classified as negative, FP indicates false positives that are incorrectly classified as positive, and TN stands for true negatives that are accurately identified as negative.
+Based on these definitions, we proceed to calculate the True Positive Rate (TPR or Recall), False Positive Rate (FPR), and Precision using the following formulas:
+TABLE III: The statistics of the four categories
+Actual Predicted Positive Negative
+Positive TP FN
+Negative FP TN
+[…]
+$\text{FPR}=\frac{FP}{FP+TN}$ (20)
+$\text{Precision}=\frac{TP}{TP+FP}$ (21)
+By calculating the FPR, TPR, and precision across different thresholds, we generate the receiver operating characteristic (ROC) curve, plotting FPR on the X-axis and TPR on the Y-axis.
+Similarly, the precision-recall (PR) curve is created, with recall on the X-axis and precision on the Y-axis.
+The evaluation metrics for DFDRNN include the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
+III-D DFDRNN’s performance in cross-validation
+We utilize 10-fold cross-validation (10-CV) to evaluate the model’s performance.
+The dataset is randomly divided into ten distinct subsets, each comprising 10% known drug-disease associations and 10% unknown pairs.
+In each iteration, one subset is used as the test set, while the other nine serve as the training set.
+This process is repeated ten times with varying random seeds, allowing for the calculation of the average and standard deviation of the performance metrics.
+The experimental outcomes, including AUROC and AUPR scores, are summarized in Table II, while the corresponding ROC and PR curves are provided in Supplementary Figures S1-S4.
+In Table II, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets.
+This performance surpassed the second-ranked NCH-DDA model by 0.7% and 1.8% in AUROC and AUPR, respectively.
+DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations.
+It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
+III-E Effect of the number of nearest neighbors
 Figure 2: (a)Line plot of AUROC as a function of top-$t$ on four benchmark datasets; (b)Line plot of AUPR as a function of top-$t$ on four benchmark datasets; (c)Line plot of the average AUROC as a function of top-$t$ on four datasets; (d)Line plot of the average AUPR as a function of top-$t$ on four datasets.
-In the context of DFDRNN, the value of top-$t$ determines the number of neighbors features aggregated during IntraDDFE, which influences the model’s performance.
+In the context of DFDRNN, the value of top-$t$ determines the number of neighbors features aggregated during SDDFE, which influences the model’s performance.
 A small value of $t$ limits the neighbor information obtained, thereby affecting the model’s performance.
 On the other hand, as $t$ increases, more neighbors are obtained, but this also introduces noise information, leading to the propagation of unnecessary data.
-The performance of DFDRNN under 10-fold cross-validation varies with $t$ on the four datasets, as demonstrated in Figure 2 for $t\in\{1,2,3,\ldots,20\}$.
+The performance of DFDRNN under 10-CV varies with $t$ on the four datasets, as demonstrated in Figure 2 for $t\in\{1,2,3,\ldots,20\}$.
 Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$.
-However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention mechanism’s aggregation capability.
+However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, demonstrating notable stability.
+This stability can be attributed to the mining capability of the SAM.
 Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across the four datasets.
 For detailed numerical values, please refer to Supplementary Tables S1 and S2.
-3.6 Ablation analysis
-3.6.1 Encoder
+III-F Ablation analysis
+III-F1 Encoder
 The encoder of DFDRNN incorporates SAM to encode drugs (or diseases) into the similarity feature (SF) and the association feature (AF).
-We introduce two dual-feature extraction modules, IntraDDFE and InterDDFE, to precisely encode drugs and diseases, with InterDDFE changing the feature (CF) of drugs (or diseases).
+We introduce two dual-feature extraction modules, SDDFE and CDDFE, to precisely encode drugs and diseases, with CDDFE changing the feature (CF) of drugs (or diseases).
 We developed various model variants to underscore these elements’ importance in the DFDRNN encoder and evaluated their performance.
@@
   • DFDRNN-AF: Remove the similarity feature and encode drugs and diseases solely as the association feature.
-  • DFDRNN-GCN: SAM was replaced by a GCN to observe the impact on the model’s performance.
-  • DFDRNN-noCF: Remove the change of features in InterDDFE, meaning that the similarity feature and the association feature will not be changed into each other after InterDDFE.
-Table 4: The AUROCs and AUPRs of DFDRNN and its four variants under 10-fold cross-validation
-Variants Fdataset Cdataset Ldataset LRSSL
+  • DFDRNN-GCN: Replace the SAM with a GCN to evaluate its effect on the model’s performance.
+  • DFDRNN-noCF: Remove the change of features in CDDFE, meaning that the similarity feature and the association feature will not be changed into each other after CDDFE.
+Table IV presents the AUROC and AUPR values of DFDRNN and its four variants under 10-CV.
+By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Gdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets.
+However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature.
+Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance.
+Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information.
+Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Gdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the CDDFE module.
+TABLE IV: The performance of DFDRNN and its four variants
+Variants Gdataset Cdataset Ldataset LRSSL
 AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
@@
 DFDRNN 0.960 0.622 0.973 0.689 0.885 0.572 0.965 0.504
-Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-fold cross-validation.
-By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Fdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets.
-However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature.
-Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance.
-Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information.
-Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Fdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the InterDDFE module.
-3.6.2 Decoder
-Figure 3: non-cross dual-domain decoder
+III-F2 Decoder
 To demonstrate the effectiveness of the cross-dual-domain decoder, an experiment was designed to compare the AUROC value and the AUPR value obtained by DFDRNN with those obtained by five variants of DFDRNN, labeled as $s_{r},s_{d},s^{non},s^{non}_{r}$, and $s^{non}_{d}$.
@@
 The specific implementation of the non-cross dual-domain decoder is described as follows.
+Figure 3: non-cross dual-domain decoder
 ${s^{non}_{r}}$ $=\mathbf{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
@@
 ${s^{non}}$ $={(s^{non}_{r}+(^{non}_{d})^{T})/2.0}$ (24)
-The heatmaps, a crucial visual representation of the results, are shown in Figure 4 across the four datasets: Fdataset, Cdataset, Ldataset, and LRSSL.
+The heatmaps, a crucial visual representation of the results, are shown in Figure 4 across the four datasets: Gdataset, Cdataset, Ldataset, and LRSSL.
 DFDRNN’s consistent outperformance in achieving the highest AUROC and AUPR values across all four datasets underscores the significance of the cross-dual-domain decoder.
@@
 Figure 4: Performance of different decoding methods: (a) AUROC heatmap; (b) AUPR heatmap.
-3.7 Identification of candidates for new diseases
+III-G Identification of candidates for new diseases
 In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment.
-Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic (Singh et al., 2020).
+Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic [39].
 Finding suitable drugs for emerging diseases is a complex and crucial task.
-To evaluate DFDRNN’s predictive ability in identifying potential candidate drugs for new diseases, we conducted leave-one-out cross-validation (LOOCV) on the Fdataset dataset.
-Specifically, for each disease $d_{x}$, all drug-disease associations related to $d_{x}$ were excluded from the analysis and treated as test samples.
-The remaining drug-disease associations were used as training samples, relying solely on the similarity between disease $d_{x}$ and other diseases to predict its candidate drugs.
-Figure 5: The performance of all methods in predicting potential drugs for new diseases on Fdataset: (a) Receiver operating characteristic (ROC) curves of prediction results obtained by applying DFDRNN and other competitive methods; (b) precision-recall (PR) curves of prediction results obtained by applying DFDRNN and other competitive methods.
-In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated.
-The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCH-DDA by 0.99%.
+To assess DFDRNN’s ability to predict candidate drugs for new diseases, we applied leave-one-out cross-validation (LOOCV) using the Gdataset.
+For each disease $d_{x}$, all drug-disease associations linked to $d_{x}$ were removed and treated as test data, while the remaining associations served as training data.
+The prediction relied on the similarity between $d_{x}$ and other diseases to identify its candidate drugs.
+Figure 5: The performance of all approaches in predicting candidate drugs for new diseases on Gdataset: (a) ROC curves illustrating the prediction outcomes of DFDRNN alongside other competitive models; (b) PR curves illustrating the prediction outcomes of DFDRNN alongside other competitive models.
+In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Gdataset is illustrated.
+The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCH-DDA by 0.81%.
 Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
-3.8 Case study
-To assess the real-world performance of DFDRNN, we conducted case studies on two diseases (Alzheimer’s disease (AD) and Parkinson’s disease (PD)) from the Fdataset.
-We utilized all drug-disease associations in the Fdataset as the training set to train DFDRNN and predict the association probabilities for all unknown drug-disease pairs.
-Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD.
-To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB (Knox et al., 2024), PubChem (Kim et al., 2016), CTD (Davis et al., 2023), DrugCentral (Avram et al., 2021), and ClinicalTrials.
+III-H Case study
+To evaluate the practical effectiveness of DFDRNN, we performed case studies on two diseases, Alzheimer’s disease (AD) and Parkinson’s disease (PD), from the Gdataset.
+Using all drug-disease associations within the Gdataset as the training data, we trained DFDRNN to predict association probabilities for all unknown drug-disease pairs.
+We then ranked the drugs in descending order of predicted probabilities, identifying the top 10 candidate drugs for AD and PD.
+To validate DFDRNN’s predictions, we referenced reliable sources and clinical trials, including DrugBank [25], PubChem [40], CTD [28], DrugCentral [41], and ClinicalTrials.
 AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder.
-Table 5 displays the top 10 drugs predicted by DFDRNN for AD.
+Table V displays the top 10 drugs predicted by DFDRNN for AD.
 Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD.
@@
 It can also be utilized for the prevention or symptomatic treatment of influenza A.
-Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment (Erkulwater and Pillai, 1989).
+Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment [42].
 Carbamazepine functions as an anticonvulsant and analgesic drug.
-Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients (Lemke, 1995).
+Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients [43].
 DFDRNN predicts an association with carbamazepine, which has been validated by CTD.
-Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
-Table 5: The top 10 DFDRNN-predicted candidate drugs for AD
+Moreover, all the other drugs listed in Table V have been confirmed by relevant authoritative data sources, with a success rate of 100%.
+TABLE V: The top 10 candidate drugs for AD are ranked based on the DFDRNN prediction score
 Rank Drug ID Drug Name Evidence
@@
 10 DB01262 Decitabine CTD
-Table 6: The top 10 DFDRNN-predicted candidate drugs for PD
+TABLE VI: The top 10 candidate drugs for PD are ranked based on the DFDRNN prediction score
 Rank Drug ID Drug Name Evidence
@@
 Parkinson’s disease (PD) is the second most common neurogenic disorder after AD.
-Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN.
+Table VI presents the top 10 potential candidate drugs for PD predicted by DFDRNN.
 Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources.
@@
 The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
-4 Conclusion
+IV Conclusion
 Our study introduces DFDRNN, a neural network model designed for drug repositioning.
-DFDRNN uses IntraDDFE and InterDDFE modules to extract the similarity and association features of drugs (or diseases), respectively, achieving precise encoding of drugs (or diseases).
-Compared with the IntraDDFE module, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature.
-In both modules, we use the self-attention mechanism to aggregate the neighbor information of drugs (and diseases).
+DFDRNN uses the SDDFE and CDDFE modules to extract the similarity and association features for drugs (or diseases), respectively, ensuring more appropriate encoding of drugs (or diseases).
+Compared with the SDDFE module, the CDDFE module changes the similarity (or association) feature into the association (or similarity) feature.
+In both modules, we use the SAM to mining the neighbor information of drugs (and diseases).
 Additionally, the DFDRNN model fuses three similarity (or association) features into a new similarity (or association) feature and uses a layer attention mechanism to calculate the final similarity (or association) feature.
 Finally, a cross-dual-domain decoder process is adopted to predict drug-disease associations in both the drug and disease domains.
-DFDRNN performs best in 10-fold cross-validation across four benchmark datasets and achieves the highest AUROC in leave-one-out cross-validation.
-The experimental results confirm the effectiveness of our proposed DFDRNN model in drug repositioning.
+DFDRNN performs best in 10-CV across four benchmark datasets and achieves the highest AUROC in LOOCV.
+The experimental results confirm the effectiveness of our proposed DFDRNN in drug repositioning.
 For future work, we would like to investigate whether specific enhancements, such as integrating temporal data and handling multi-modal data, could further enhance the performance of DFDRNN.
@@
 We aim to identify and implement the appropriate techniques and models for these enhancements.
-Furthermore, we intend to explore the integration of additional biological entities, such as miRNA and proteins, for drug repositioning.
+Furthermore, we plan to investigate incorporating other biological entities, such as miRNAs and proteins, for drug repositioning.
 Acknowledgments

### v4

@@
 Case studies on two diseases show that the proposed DFDRNN model can be applied in real-world scenarios, demonstrating its significant potential in drug repositioning.
-Index Terms:
-Drug Repositioning, Self-Attention Mechanism, Dual-Feature Extraction, Cross-Dual-Domain
-I Introduction
+keywords:
+American Chemical Society,
+[…]
+IR,NMR,UV
+1 Introduction
 Drug development is a significant yet complex and time-consuming process.
 It consists of three main stages: the discovery stage, the preclinical stage, and the clinical stage.
-Typically, the entire process takes about 10 to 15 years and costs around \$2.8 billion on average [1, 2].
+Typically, the entire process takes about 10 to 15 years and costs around \$2.8 billion on average 1, 2.
 Despite ongoing efforts and innovations, the traditional drug development model faces considerable risks, often resulting in project terminations due to concerns about drug efficacy, safety, or commercial viability.
@@
 Consequently, finding a more efficient and precise method for drug development is critically important and valuable.
-Given the high cost of drug development, computational drug repositioning has become a powerful strategy for discovering alternative uses of already approved drugs while considering their potential effects [3].
+Given the high cost of drug development, computational drug repositioning has become a powerful strategy for discovering alternative uses of already approved drugs while considering their potential effects 3.
 This approach utilizes computational and bioinformatics tools, providing greater efficiency and cost-effectiveness compared to traditional drug research and development methods.
@@
 Machine learning has been commonly used in computational drug repositioning by information extraction from biomedical data.
-Techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose [4].
-For instance, [5] introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets.
-[6] proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier.
-[7] employed random forests in their final prediction step, while other techniques such as decision trees [8], matrix factorization[9, 10], and logistic regression [11] have also been developed.
-In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by [12, 13, 14, 15].
-Unlike traditional machine learning methods, deep learning techniques leverage deep architectures to effectively capture high-level latent representations without requiring manual feature selection and tuning [16].
-Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building [17, 18, 19, 20, 21].
-For instance, [17] performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases but did not account for the differences in encoding from different network spaces.
-[18] distinguished intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases.
-[21] extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features.
+Techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose 4.
+For instance, 5 introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets.
+6 proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier.
+7 employed random forests in their final prediction step, while other techniques such as decision trees 8, matrix factorization9, 10, 11, and logistic regression 12 have also been developed.
+In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by 13, 14, 15, 16.
+Unlike traditional machine learning methods, deep learning techniques leverage deep architectures to effectively capture high-level latent representations without requiring manual feature selection and tuning 17.
+Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building 18, 19, 20, 21, 22.
+For instance, 18 performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases but did not account for the differences in encoding from different network spaces.
+19 distinguished intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases.
+22 extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features.
 However, these studies did not consider calculating information weights from different neighbors.
-Authors in [22, 23] used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding.
+Authors in 23, 24, 25 used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding.
 Most deep learning methods have primarily focused on extracting valuable features from neighborhoods.
@@
 In addition, case studies on Alzheimer’s and Parkinson’s diseases validated the reliability of DFDRNN in practical applications.
-The rest of this paper is structured as follows: Section II introduces the datasets and methods, providing details on the DFDRNN implementation; Section III explains the experimental setup and discusses the results; Section IV concludes the study and suggests future research directions.
-II Materials and Methods
-II-A Datasets
+The rest of this paper is structured as follows: Section 2 introduces the datasets and methods, providing details on the DFDRNN implementation; Section 3 explains the experimental setup and discusses the results; Section 4 concludes the study and suggests future research directions.
+2 Materials and Methods
+2.1 Datasets
 We performed experiments on four benchmark datasets.
-The first, Gdataset [24], comprises 593 drugs from DrugBank [25] and 313 diseases from the Online Mendelian Inheritance in Man (OMIM) [26], with a total of 1933 drug-disease associations.
-The second dataset, Cdataset [27], consists of 663 drugs listed in DrugBank, 409 diseases from OMIM, and 2532 known associations.
-Ldataset [17], the third dataset, contains 18,416 associations between 269 drugs and 598 diseases, derived from the Comparative Toxicogenomics Database (CTD) [28].
-Lastly, LRSSL [29] contains 763 drugs, 681 diseases, and 3051 validated drug-disease associations.
-A summary of the statistics for these datasets can be found in Table I.
-This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation [30].
-For disease similarity, it is determined based on disease phenotypes and computed via MimMiner [31], which measures similarity by analyzing the MeSH terms in the medical descriptions of diseases found in the OMIM database [32].
-TABLE I: Summary statistics of the four datasets
+The first, Gdataset 26, comprises 593 drugs from DrugBank 27 and 313 diseases from the Online Mendelian Inheritance in Man (OMIM) 28, with a total of 1933 drug-disease associations.
+The second dataset, Cdataset 29, consists of 663 drugs listed in DrugBank, 409 diseases from OMIM, and 2532 known associations.
+Ldataset 18, the third dataset, contains 18,416 associations between 269 drugs and 598 diseases, derived from the Comparative Toxicogenomics Database (CTD) 30.
+Lastly, LRSSL 31 contains 763 drugs, 681 diseases, and 3051 validated drug-disease associations.
+A summary of the statistics for these datasets can be found in Table 1.
+This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation 32.
+For disease similarity, it is determined based on disease phenotypes and computed via MimMiner 33, which measures similarity by analyzing the MeSH terms in the medical descriptions of diseases found in the OMIM database 34.
+Table 1: Summary statistics of the four datasets
 Datasets No. of drugs No. of disease No. of assocations
@@
 LRSSL 763 681 3051
-II-B Model structure
+2.2 Model structure
 Our research is focused on predicting potential associations between drugs and diseases.
@@
 The decoder then calculates drug-disease association scores separately for each domain.
-II-B1 Three key networks
+2.2.1 Three key networks
 To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity between the two drugs incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge.
@@
 $\ \qquad\qquad E_{s}=\begin{bmatrix}R&0\\ 0&D\end{bmatrix}\qquad E_{a}=\begin{bmatrix}0&A\\ A^{T}&0\end{bmatrix}$ (2)
-II-B2 Self-attention mechanism
+2.2.2 Self-attention mechanism
 The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector.
-The output of the SAM is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ [33].
+The output of the SAM is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ 35.
 In our SDDFE and CDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases.
@@
 We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer.
-Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU [34] activation function.
+Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU 36 activation function.
 Then,
@@
 $=\begin{bmatrix}\sigma(H_{r}w_{r}+b_{r})\\ \sigma(H_{d}w_{d}+b_{d})\end{bmatrix}$
-II-B3 Encoder
+2.2.3 Encoder
 The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: SDDFE module based on network $E_{s}$ and CDDFE module based on network $E_{a}$.
@@
 $\begin{bmatrix}H^{r}_{s}\\ H^{d}_{s}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{s}),\qquad\begin{bmatrix}H^{r}_{a}\\ H^{d}_{a}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{a})$ (14)
-II-B4 Decoder
+2.2.4 Decoder
 A cross-dual-domain decoder is utilized to predict drug-disease association.
@@
 3:  Initialize the weight matrix ($w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$).
-4:   for $c=1$ to $C$
+4:  for $c=1$ to $C$ do
 5:           Compute the similarity feature $H^{0}_{s}$ and the association feature $H^{0}_{a}$ with Equation (9).
-6:        for $\ell=$ to $L-1$
-7:              Compute SDDFE $\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a}$ with Equation (10).
-8:              Compute CDDFE $\tilde{H}^{\ell+1}_{s},\tilde{H}^{\ell+1}_{a}$ with Equation (11).
-9:              Fuse the same feature and the input feature yields $H^{\ell+1}_{s}$ and $H^{\ell+1}_{a}$ with Equation (12) (13)
-10:        end for
-11:        Utilizing layer attention mechanism to compute the final embedding with Equation (14)
-12:        Obtain Computer the prediction score matrix $s$ with Equation (15)(16)(17)
-13:        Update the parameters $w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$ with Equation (18).
-14:        $c\leftarrow c+1$.
-15:   end for
-II-C Optimization
+6:
+7:     for $\ell=$ to $L-1$ do
+8:                    Compute SDDFE $\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a}$ with Equation (10).
+9:                    Compute CDDFE $\tilde{H}^{\ell+1}_{s},\tilde{H}^{\ell+1}_{a}$ with Equation (11).
+10:                    Fusing the same feature with Equations (12) and (13) produces $H^{\ell+1}_{s}$ and $H^{\ell+1}_{a}$.
+11:
+12:     end for
+13:           Final embedding via layer attention and Equation (14).
+14:           Obtain Computer the prediction score matrix $s$ with Equation (15)(16)(17)
+15:           Update the parameters $w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$ with Equation (18).
+16:           $c\leftarrow c+1$.
+17:  end for
+2.3 Optimization
 Predicting drug-disease associations can be treated as a binary classification problem, where the goal is to determine if a specific drug is associated with a specific disease.
@@
 $loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{ij})+\sum_{(i,j)\in y^{-}}\log(1-s_{ij}))$ (18)
-The DFDRNN model is optimized using the Adam optimizer [35], with parameters initialized through Xavier initialization [36].
-To enhance generalization, we randomly drop edges and apply regular dropout techniques [37].
-III Experiments and Results
-III-A Baseline methods
+The DFDRNN model is optimized using the Adam optimizer 37, with parameters initialized through Xavier initialization 38.
+To enhance generalization, we randomly drop edges and apply regular dropout techniques 39.
+3 Experiments and Results
+3.1 Baseline methods
 We compared DFDRNN against six state-of-the-art methods for predicting drug-disease associations across the above four datasets. These methods include:
-$\bullet$ LAGCN [17]: This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
-$\bullet$ DRHGCN [18]: This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
-$\bullet$ DRWBNCF [38]: This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
-$\bullet$ NCH-DDA [21]: This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases.
+$\bullet$ LAGCN 18: This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
+$\bullet$ DRHGCN 19: This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
+$\bullet$ DRWBNCF 40: This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
+$\bullet$ NCH-DDA 22: This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases.
 Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
-$\bullet$ HDGAT [23]: This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
-$\bullet$ DRGBCN [20]: This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
-III-B Experimental parameter configuration
+$\bullet$ HDGAT 24: This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
+$\bullet$ DRGBCN 21: This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
+3.2 Experimental parameter configuration
 The model’s parameter values are as follows: DFDRNN’s feature embedding dimension is 128; the number of network layers is 3; the number of attention heads is p = 2; the dropout rate is 0.4; the edge dropout rate is 0.5 on the Ldataset and 0.2 on other datasets; the learning rate is 0.008; and the maximum training epoch is 800.
 Additionally, we set top-$t$ to 7 based on parameter sensitivity experiments.
-TABLE II: AUROC and AUPR results for DFDRNN and comparative models using 10-times repeated 10-fold cross-validation
+Table 2: AUROC and AUPR results for DFDRNN and comparative models using 10-times repeated 10-fold cross-validation
 Metrics Model Gdataset Cdataset Ldataset LRSSL Avg.
@@
   • The highest values in each column are highlighted in bold, while the second-highest values are underlined
-III-C Evaluation metrics
+3.3 Evaluation metrics
 We classify the prediction outcomes into four categories.
-As illustrated in Table III, TP denotes true positives that are correctly identified as positive, FN represents false negatives that are wrongly classified as negative, FP indicates false positives that are incorrectly classified as positive, and TN stands for true negatives that are accurately identified as negative.
+As illustrated in Table 3, TP denotes true positives that are correctly identified as positive, FN represents false negatives that are wrongly classified as negative, FP indicates false positives that are incorrectly classified as positive, and TN stands for true negatives that are accurately identified as negative.
 Based on these definitions, we proceed to calculate the True Positive Rate (TPR or Recall), False Positive Rate (FPR), and Precision using the following formulas:
-TABLE III: The statistics of the four categories
+Table 3: The statistics of the four categories
 Actual Predicted Positive Negative
@@
 The evaluation metrics for DFDRNN include the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
-III-D DFDRNN’s performance in cross-validation
+3.4 DFDRNN’s performance in cross-validation
 We utilize 10-fold cross-validation (10-CV) to evaluate the model’s performance.
@@
 This process is repeated ten times with varying random seeds, allowing for the calculation of the average and standard deviation of the performance metrics.
-The experimental outcomes, including AUROC and AUPR scores, are summarized in Table II, while the corresponding ROC and PR curves are provided in Supplementary Figures S1-S4.
-In Table II, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets.
+The experimental outcomes, including AUROC and AUPR scores, are summarized in Table 2, while the corresponding ROC and PR curves are provided in Supplementary Figures S1-S4.
+In Table 2, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets.
 This performance surpassed the second-ranked NCH-DDA model by 0.7% and 1.8% in AUROC and AUPR, respectively.
@@
 It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
-III-E Effect of the number of nearest neighbors
+3.5 Effect of the number of nearest neighbors
 Figure 2: (a)Line plot of AUROC as a function of top-$t$ on four benchmark datasets; (b)Line plot of AUPR as a function of top-$t$ on four benchmark datasets; (c)Line plot of the average AUROC as a function of top-$t$ on four datasets; (d)Line plot of the average AUPR as a function of top-$t$ on four datasets.
@@
 For detailed numerical values, please refer to Supplementary Tables S1 and S2.
-III-F Ablation analysis
-III-F1 Encoder
+3.6 Ablation analysis
+3.6.1 Encoder
 The encoder of DFDRNN incorporates SAM to encode drugs (or diseases) into the similarity feature (SF) and the association feature (AF).
@@
   • DFDRNN-noCF: Remove the change of features in CDDFE, meaning that the similarity feature and the association feature will not be changed into each other after CDDFE.
-Table IV presents the AUROC and AUPR values of DFDRNN and its four variants under 10-CV.
+Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-CV.
 By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Gdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets.
@@
 Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Gdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the CDDFE module.
-TABLE IV: The performance of DFDRNN and its four variants
+Table 4: The performance of DFDRNN and its four variants
 Variants Gdataset Cdataset Ldataset LRSSL
@@
 DFDRNN 0.960 0.622 0.973 0.689 0.885 0.572 0.965 0.504
-III-F2 Decoder
+3.6.2 Decoder
 To demonstrate the effectiveness of the cross-dual-domain decoder, an experiment was designed to compare the AUROC value and the AUPR value obtained by DFDRNN with those obtained by five variants of DFDRNN, labeled as $s_{r},s_{d},s^{non},s^{non}_{r}$, and $s^{non}_{d}$.
@@
 Figure 4: Performance of different decoding methods: (a) AUROC heatmap; (b) AUPR heatmap.
-III-G Identification of candidates for new diseases
+3.7 Identification of candidates for new diseases
 In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment.
-Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic [39].
+Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic 41.
 Finding suitable drugs for emerging diseases is a complex and crucial task.
@@
 Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
-III-H Case study
+3.8 Case study
 To evaluate the practical effectiveness of DFDRNN, we performed case studies on two diseases, Alzheimer’s disease (AD) and Parkinson’s disease (PD), from the Gdataset.
@@
 We then ranked the drugs in descending order of predicted probabilities, identifying the top 10 candidate drugs for AD and PD.
-To validate DFDRNN’s predictions, we referenced reliable sources and clinical trials, including DrugBank [25], PubChem [40], CTD [28], DrugCentral [41], and ClinicalTrials.
+To validate DFDRNN’s predictions, we referenced reliable sources and clinical trials, including DrugBank 27, PubChem 42, CTD 30, DrugCentral 43, and ClinicalTrials.
 AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder.
-Table V displays the top 10 drugs predicted by DFDRNN for AD.
+Table 5 displays the top 10 drugs predicted by DFDRNN for AD.
 Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD.
@@
 It can also be utilized for the prevention or symptomatic treatment of influenza A.
-Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment [42].
+Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment 44.
 Carbamazepine functions as an anticonvulsant and analgesic drug.
-Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients [43].
+Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients 45.
 DFDRNN predicts an association with carbamazepine, which has been validated by CTD.
-Moreover, all the other drugs listed in Table V have been confirmed by relevant authoritative data sources, with a success rate of 100%.
-TABLE V: The top 10 candidate drugs for AD are ranked based on the DFDRNN prediction score
+Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
+Table 5: The top 10 candidate drugs for AD are ranked based on the DFDRNN prediction score
 Rank Drug ID Drug Name Evidence
@@
 10 DB01262 Decitabine CTD
-TABLE VI: The top 10 candidate drugs for PD are ranked based on the DFDRNN prediction score
+Table 6: The top 10 candidate drugs for PD are ranked based on the DFDRNN prediction score
 Rank Drug ID Drug Name Evidence
@@
 Parkinson’s disease (PD) is the second most common neurogenic disorder after AD.
-Table VI presents the top 10 potential candidate drugs for PD predicted by DFDRNN.
+Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN.
 Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources.
@@
 The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
-IV Conclusion
+4 Conclusion
 Our study introduces DFDRNN, a neural network model designed for drug repositioning.
@@
 Furthermore, we plan to investigate incorporating other biological entities, such as miRNAs and proteins, for drug repositioning.
-Acknowledgments
-This work was supported by the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China under grants 62172072 and 62272115.
+{acknowledgement}
+The authors thank the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China (Grants 62172072 and 62272115) for their financial support.
+5 Data Availability
+The implementation and data of the DFDRNN model are available at https://github.com/siekersun/DFDRNN.git.
+{suppinfo}
+  • manuscript.pdf: Manuscript PDF File
+  • manuscript.zip: This zip package contains files for an academic paper written in LaTeX, including the main document, references, and document class files for formatting.
+  • supplementary.docx:Supporting Information for Publication